### v1

# Learning to Solve Multi-Objective Routing Problems on Multigraphs

Abstract
Learning-based methods for routing have gained significant attention in recent years, both in single-objective and multi-objective contexts. However, the multigraph setting, where multiple paths with distinct attributes can exist between destinations, has largely been overlooked, despite its high practical relevancy. In this paper, we introduce two neural approaches to address multi-objective routing on multigraphs. Our first approach works directly on the multigraph, by autoregressively selecting edges until a tour is completed. On the other hand, our second model first prunes the multigraph into a simple graph and then builds routes. We validate both models experimentally and find that they demonstrate strong performance across a variety of problems, including the Traveling Salesman Problem (TSP) and Capacitated Vehicle Routing Problem (CVRP)111Our code will be made publicly available upon acceptance of the paper..
1 Introduction
Routing problems, including the well-known Traveling Salesman Problem (TSP), are central for optimizing operations in logistics, delivery and beyond. However, these problems are generally NP-hard and difficult to solve exactly in practical, real-time settings. While exact solvers provide high-quality solutions, they do not scale well. At the same time, hand-crafted heuristics are often efficient but time-consuming to design and typically lack generalizability across problem variants.
Recently, neural solvers have emerged as a promising alternative, offering speed, accuracy, and broad applicability across routing problems (Kool, van Hoof, and Welling 2018; Ye et al. 2024; Drakulic, Michel, and Andreoli 2025). However, most learning-based approaches focus on Euclidean settings, where customers are points in a plane and travel costs are based on the Euclidean distance. Real-world road networks are far more complex and travel cost is usually asymmetric, affected by elevation, one-way streets and traffic. Further, when there are multiple conflicting criteria of importance, there can be multiple paths between customers that cannot be chosen between apriori.
In such cases, a directed multigraph, allowing multiple, distinct paths between the same locations, is a more realistic and powerful input-structure. In fact, using this representation has empirically been shown to increase solution quality significantly compared to the simple graph representation (Ben Ticha et al. 2017). Yet this comes at the cost of increased complexity, underscoring the need for new methods capable of handling multigraph-based routing efficiently.
The challenge intensifies in Multi-Objective (MO) routing problems, such as the Multi-Objective TSP (MOTSP), where several criteria (e.g., distance and time) must be optimized simultaneously. Even without multigraphs, MO problems are hard: not only because their single-objective counter-part is often NP-hard, but also because the number of Pareto optimal solutions can increase exponentially with the size of the problem (Ehrgott and Gandibleux 2000).
Pareto Set Learning (PSL) has recently emerged as a powerful framework for MO routing. By training offline on similar instances, PSL approximates the Pareto set quickly during inference. Moreover, since the framework is based on reinforcement learning, there is no need for oracle solvers for supervised training. These methods have shown strong results in Euclidean settings (Lin, Yang, and Zhang 2022; Fan et al. 2025), but there is currently no general PSL framework for routing on multigraphs or in non-Euclidean settings, despite their higher practical relevance.
We attempt to bridge this gap. Concretely, our main contributions are:
We extend PSL to handle multigraph inputs, where the model selects not only the next node but also the path taken. To our knowledge, this is the first neural solver designed for such input structures. Although we treat multi-objective optimization, we hypothesize that our approach can be useful in single-objective constrained and stochastic cases too.
As a byproduct, our model constitutes the first general-purpose, multi-objective solver capable of handling a variety of routing problems on asymmetric graphs.
In order to do this, we propose two separate models: One edge-based that is simple and accurate but relatively slow and one dual-task model that first sparsifies the multigraph into a simple graph and then selects routes. The latter is faster but more complex.
Experimentally, we show that our approach achieves competitive results for several MO routing problems on asymmetric graphs and multigraphs, including the MOTSP, Multi-Objective Capacitated Vehicle Routing Problem (MOCVRP) and MOTSP with time-windows.
2 Related Work
2.1 Learning for Routing
Motivated by the success of end-to-end learning rather than hand-crafted heuristics in fields such as image classification and language modeling, neural heuristics for routing have gained increasing attention during the last years. Broadly, methods can be classified by how they obtain the tours. Two prominent alternatives are autoregressive construction, based on iterative building of tours one node at a time, and non-autoregressive construction, that outputs heatmaps in a one-shot fashion for down-stream decoding with some search method.
Early works in the first category include Vinyals, Fortunato, and Jaitly (2015), which utilized an oracle TSP solver for supervised learning, and Bello et al. (2017), which introduced reinforcement learning for training. By using attention models and inherent symmetries in the problem, first Kool, van Hoof, and Welling (2018) and then Kwon et al. (2020) obtained close to 0 optimality gap on various routing problems. Regarding asymmetric problems, Kwon et al. (2021) proposed MatNet, a mixed-attention architecture working on bi-partite graphs. More recent works are Jin et al. (2023); Drakulic et al. (2023); Chen et al. (2024); Bi et al. (2024). These papers have improved results, can handle constraints better and can scale to even larger problems than previously. Another notable example is Drakulic, Michel, and Andreoli (2025), where one model learns to solve a multitude of combinatorial optimization problems, including examples from routing.
In the second category, some notable works are Joshi, Laurent, and Bresson (2019); Fu, Qiu, and Zha (2021); Qiu, Sun, and Yang (2022); Sun and Yang (2023). Typically, these methods perform worse than autoregressive approaches, but scale better to large instances due to lower space and time complexity. An interesting recent work combining the two approaches to solve large instances (up to nodes for the TSP) is Ye et al. (2024).
2.2 Multi-Objective Learning for Routing
The literature on multi-objective routing is more sparse than in the single-objective case. Some early works (Li, Zhang, and Wang 2021; Wu, Wang, and Zhang 2020) utilize multiple models, where one model is responsible for one preference between the objectives. This quickly becomes infeasible as the number of models required grows exponentially with the number of objectives (Ehrgott 2005). Pareto-set learning for multi-objective routing using a single model was proposed in Lin, Yang, and Zhang (2022). They used the hyper-networks of Navon et al. (2021) to condition the decoder on the preference. Recent papers that feature improved results with a similar conditioning method include Fan et al. (2025); Lu et al. (2024). An alternative approach is to use meta-learning to train a single model that can be quickly fine-tuned over a few steps to solve individual subproblems. Two representative works in this direction are Zhang et al. (2021); Chen et al. (2023a).
Regarding asymmetric problems in the context of multi-objective optimization, the only existing learning-based methods we are aware of are those proposed in Santiyuda et al. (2024) and Zhou, Kong, and Wang (2025). Notably, both of these methods are tailored to particular problem types and evaluated within narrowly defined domains. In contrast, our approach is designed with a very general framework, capable of handling a broad class of routing problems.
2.3 Routing on Multigraphs
In the multi-objective case, a multigraph is often a more suitable representation of a road network than a simple graph. In fact, this network model has become more and more common for MO shortest path problems. One solution is genetic algorithms, see for instance Beke, Weiszer, and Chen (2021) for a comprehensive benchmarking study. Other studies have proposed A∗-related algorithms, e.g., AMOA∗ (Weiszer, Burke, and Chen 2020) and MOMGA∗ (Liu, Chen, and Weiszer 2022), to efficiently obtain a Pareto set of shortest paths.
The multigraph framework has also been applied to routing problems, but to a lower extent. The first study in this regard was Garaix et al. (2010), who designed a simple heuristic and branch-and-price method for a single-objective on-demand transportation problem. Other works proposing either exact solvers or heuristics for single-objective constrained problems or multi-objective problems are Lai, Caliskan Demirag, and Leung (2016); Ben Ticha et al. (2019); Tikani, Setak, and Demir (2021). In addition, a particularly interesting study is Ben Ticha et al. (2017), where it is clearly concluded that a multigraph setup yields significantly lower costs compared to a simple graph representation for the capacitated vehicle routing problem with time-windows. On real-world instances, a solution cost reduction of up to is reported. At the same time, the computation time increases significantly under this perspective, and the authors highlight the need for new heuristics.
We believe learning-based methods can fill this niche, as they have proven to be efficient, effective and broadly applicable. To the best of our knowledge, this work presents the first learning-based routing approach to address problems on multigraphs.
3 Multi-Objective Optimization
A general multi-objective optimization problem can be written as
where and . Unlike the Single-Objective (SO) case, the solution to such a problem is not one objective value with corresponding optimal solutions but a set of values known as the Pareto Front (PF) with corresponding solutions in the Pareto Set (PS). Formally, we define the PS as the set
where denotes that dominates , meaning
Correspondingly, the Pareto Front (PF) is defined as the image of the Pareto Set under , that is,
Intuitively, the PS represents solutions for which no strictly better alternative exists.
To obtain this set, a typical method is decomposition through scalarization. The idea is to translate the MO problem into several SO problems that can be solved in parallel or consecutively. Let such that and be a vector that controls the preference among the objectives. Then, the most straight-forward scalarization is the Linear Scalarization (LS) . The LS subproblem is then , which is a standard SO problem. Unfortunately, it is well-known that there may be solutions in the Pareto Set that do not solve an LS subproblem for any preference (Ehrgott 2005). The Chebyshev Scalarization has more favorable properties. Let be an ideal value for objective . Then the Chebyshev scalarized cost is
Given a solution in the PS, there always exists a preference and corresponding subproblem for which the solution is optimal (Choo and Atkins 1983).
In our context of multigraphs, it is interesting to consider how the problem simplifies under scalarization. We note that under LS, the MOTSP on a multigraph can be simplified to a simple graph by pruning edges that have sub-optimal LS cost compared to a parallel edge. Under Chebyshev scalarization or for more complicated problems this simple method will not work. Instead, we propose to learn which parallel edge is optimal while simultaneously learning to construct the optimal route.
4 GNN-Based Multigraph Solver
In the following, we detail two approaches based on Graph Neural Networks (GNNs) for solving MO multigraph routing problems. We call our method GNN-based Multigraph Solver (GMS) and propose an edge-based variant, which is slow but simple, as well as an architecture with two decoders, which is faster but more complex. Both of these utilize the Graph Edge Attention Network (GREAT) of Lischka et al. (2024) as the encoder, the reason being that it 1) Can handle input data structured as multigraphs; 2) Outputs edge-embeddings; 3) Has been applied in an autoregressive-framework and shown competitive performance for non-Euclidean SO problems. However, note that it is possible to use other architectures in the encoder, such as e.g., the anisotropic GNN of Sun and Yang (2023).
Specifically, we utilize the Node-Based (NB) version of GREAT. We refer the reader to Lischka et al. (2024) for details. Here, it suffices to say that the core component of an NB GREAT-layer is an attention sublayer, that utilizes attention scores and to compute a temporary node-feature for each node according to
Here, is the concatenation operation, denotes an edge embedding for edge , and are trainable weight matrices and , denote outgoing and incoming edges respectively. New edge embeddings are then computed according to
where and denote the start and end nodes of the directed edge . The attention layers are arranged into a single GREAT-layer together with feed-forward layers, residual connections and normalization according to the original transformer architecture (Vaswani et al. 2017). Note that the residual connections are important as they allow parallel edges to have differing embeddings.
4.1 Edge-Based GMS
We visualize Edge-Based GMS (GMS-EB) in Figure 1. The encoder, consisting of GREAT-layers, outputs edge embeddings. Using these, the decoder constructs valid tours auto-regressively. As in Lin, Yang, and Zhang (2022); Navon et al. (2021) we utilize a hyper-network to generate the weights of the decoder conditioned on the current preference. The decoder is an edge-based variant of the Multi-Pointer (MP) decoder from Jin et al. (2023), which we detail in Appendix A. Given the context of the current node in the tour, the decoder outputs probabilities of traversing edges starting in that node.
Note that the underlying graphs in routing problems are usually complete. Hence there are edges in a graph of nodes. Moreover, the decoder needs to be run times to construct a full tour and we usually roll out sample trajectories. Hence an edge-based decoder scales according to compared to a node-based that scales according to . We find that this is a bottleneck in terms of memory requirements and inference time. Hence we propose a second architecture that utilizes a node-based decoder.
4.2 Dual Head GMS
Using GREAT, we can obtain node-embeddings by not applying (2) in the final layer. The problem is that the action of selecting a node is not sufficient, as a path selection is also needed. Thus, we need to ensure that the action of choosing a node uniquely defines which edge to choose, which is achieved by pruning parallel edges leaving a unique one between each node-pair. In general, choosing this edge is not trivial. Thus, we propose to learn the choice using a dual decoder setup.
We call this approach Dual Head GMS (GMS-DH) and it is visualized in Figure 2. The key is that we insert a second decoder before the final encoder layer. This decoder uses the intermediate edge-embeddings to select which parallel edge to keep. The selection is fed back into the encoder, which only propagates the embedding of the remaining edges and converts them to node-embeddings. We also find it beneficial to add a small number of standard transformer layers after the final GREAT-layer in order to obtain more expressive node-embeddings. We detail the architecture for the selection-decoder in Appendix A. For the routing-decoder, we use a preference-conditioned version of the Multi-pointer decoder (Jin et al. 2023). When applying GMS-DH to simple graphs, we simply remove the first decoder. Also note that this architecture ensures that the first layers only need to be run once for each instance, which reduces the run-time compared to selecting edges before encoding.
From a broader perspective, our approach in GMS-DH can be viewed as a mixture of non-autoregressive and autoregressive learning. As a drawback, the edge-choice can no longer depend on the routing context, instead it must be chosen before the roll-out is started.
4.3 Training Algorithm
Our first proposed model, GMS-EB, is trained using the Multi-Objective REINFORCE from Lin, Yang, and Zhang (2022). In short, given a preference and instance , the model parameterizes a policy . A tour is associated with a reward according to (1)
The objective is to maximize
Additionally, we form a baseline by rolling out tours for the same instance and averaging the reward, according to the POMO framework (Kwon et al. 2020), to stabilize the training. The policy gradient is utilized to update the model weights and it is approximated, given one preference , according to
where is the batch size of instances and the advantage is
For the second model, the training is more complicated. We can view each head as an agent. The selection-head takes actions corresponding to which edge to choose between each node pair. It parameterizes a policy . The routing-head uses the selection, which we denote by . It parameterizes a policy and obtains a scalar reward according to (3). With a fixed , the objective for this head is to maximize
On the other hand, the task of the selection-head is to ensure that the edge selection is optimal. Let denote the optimal tour given edge selection . Then, define
As the optimal tour is not available, we approximate it using sampled tours :
A baseline for the selection-head is formed by averaging over samples while a baseline for the routing-head is formed by averaging over the samples for a fixed and . Thus, we approximate the policy gradients with
For easy reference, we outline the training of GMS-DH in Algorithm 1.
Input: Preference distribution , instance distribution , batch size , number of samples per instance head 1, , and head 2,
In addition, for both models we utilize a curriculum learning approach to speed up training. We start by training on small graph sizes and then gradually increase the problem size. For GMS-DH in the multigraph setting, we also start by training the routing-head on simple graphs to ensure that the approximation (4) is accurate even during the initial epochs of multigraph training. Further details regarding the curriculum learning can be found in Appendix B.
5 Experimental Studies
In this section we show the empirical performance for our two proposed methods on several routing problems across multiple instance sizes and distributions. We compare against evolutionary algorithms, neural baselines and state-of-the-art single-objective heuristics.
5.1 Benchmarking Setup
Problems: In the asymmetric simple graph case we consider two standard routing problems: The Multi-Objective Traveling Salesman Problem (MOTSP) and Multi-Objective Capacitated Vehicle Routing Problem (MOCVRP). Specifically, we consider bi-objective variants of these. In the MOTSP, we sample two distances for each node pair. These are summarized over the tour to yield the objective values. Regarding the distances, we consider both the TMAT (Kwon et al. 2021) and XASY distributions (Gaile et al. 2022), where the former obeys the triangle inequality and the latter does not - making it more difficult. For the MOCVRP, we sample one distance according to these distributions. We then let the first objective be given by the total distance of a tour and the second objective by the makespan, which is the longest trip for a single vehicle. This is a standard benchmark problem in literature (Lin, Yang, and Zhang 2022).
In the multigraph setting, we consider the Multigraph MOTSP (MGMOTSP) with two objectives, where we sample two-dimensional edge distances using distributions called FIX and FLEX. Here, denotes the number of edges between each node pair. In the FLEX-distribution, we sample edge distances independently and remove edges which are dominated. In the FIX distribution we first sample edges independently, but then rearrange by sorting, so that the an edge that is best in one objective is the worst in the other objective and so on. Thus, in the FIX-distribution there are always exactly edges between each node pair, whereas FLEX has at most edges.
In Appendix C, we provide more in-depth explanations of the problems and distributions. We also provide results on Euclidean problems in Appendix E.
| TMAT50 | TMAT100 | XASY50 | XASY100 | ||||||||||
| HV | Gap | Time | HV | Gap | Time | HV | Gap | Time | HV | Gap | Time | ||
| MOTSP | LKH | 0.58 | 0.00% | (6.4m) | 0.63 | 0.00% | (29m) | 0.83 | 0.00% | (6.9m) | 0.90 | 0.00% | (32m) |
| OR Tools | 0.54 | 6.13% | (29m) | 0.59 | 7.12% | (2.5h) | 0.77 | 7.25% | (24m) | 0.85 | 6.06% | (1.8h) | |
| MOEA/D | 0.53 | 9.51% | (3.7h) | 0.50 | 20.9% | (13h) | 0.73 | 12.68% | (3.6h) | 0.70 | 22.22% | (13h) | |
| MBM | 0.50 | 13.5% | (3.9s) | 0.56 | 11.2% | (19s) | 0.75 | 10.37% | (3.7s) | 0.84 | 7.41% | (19s) | |
| MBM (aug) | 0.52 | 10.2% | (27s) | 0.58 | 9.17% | (2.4m) | 0.76 | 8.28% | (27s) | 0.85 | 6.34% | (2.4m) | |
| GMS-EB | 0.57 | 1.50% | (25s) | 0.63 | 1.45% | (3.6m) | 0.81 | 2.94% | (25s) | 0.88 | 3.03% | (3.6m) | |
| GMS-DH | 0.57 | 2.36% | (4.1s) | 0.62 | 2.76% | (20s) | 0.80 | 4.12% | (3.9s) | 0.86 | 4.87% | (20s) | |
| GMS-EB (aug) | 0.57 | 1.14% | (3.3m) | 0.63 | 1.13% | (29m) | 0.82 | 2.06% | (3.3m) | 0.88 | 2.76% | (28m) | |
| GMS-DH (aug) | 0.57 | 1.76% | (28s) | 0.62 | 2.19% | (2.6m) | 0.81 | 2.84% | (28s) | 0.87 | 3.85% | (2.6m) | |
| TMAT50 | TMAT100 | XASY50 | XASY100 | ||||||||||
| MOCVRP | MOEA/D | 0.41 | 13.98% | (24h) | 0.19 | 58.70% | (76h) | 0.65 | 13.88% | (24h) | 0.42 | 47.23% | (83h) |
| MBM | 0.47 | 1.33% | (6.2s) | 0.43 | 4.50% | (27s) | 0.70 | 6.57% | (6.2s) | 0.60 | 24.47% | (27s) | |
| MBM (aug) | 0.47 | 0.99% | (41s) | 0.44 | 3.29% | (3.3m) | 0.72 | 4.13% | (42s) | 0.64 | 19.60% | (3.4m) | |
| GMS-EB | 0.47 | 0.25% | (36s) | 0.45 | 0.57% | (4.2m) | 0.74 | 1.89% | (36s) | 0.79 | 1.23% | (4.3m) | |
| GMS-DH | 0.47 | 0.76% | (6.8s) | 0.45 | 1.46% | (32s) | 0.72 | 4.13% | (6.6s) | 0.77 | 3.60% | (33s) | |
| GMS-EB (aug) | 0.47 | 0.00% | (5.0m) | 0.45 | 0.00% | (33m) | 0.75 | 0.00% | (4.9m) | 0.80 | 0.00% | (48m) | |
| GMS-DH (aug) | 0.47 | 0.36% | (47s) | 0.45 | 0.79% | (4.3m) | 0.74 | 1.65% | (47s) | 0.79 | 1.21% | (4.7m) | |
| FLEX2-50 | FLEX2-100 | FIX2-50 | FIX2-100 | ||||||||||
| MGMOTSP | LKH | 0.90 | 0.00% | (6.3m) | 0.94 | 0.00% | (31m) | 0.92 | 0.00% | (6.3m) | 0.95 | 0.00% | (30m) |
| OR Tools | 0.86 | 4.21% | (24m) | 0.91 | 3.46% | (1.8h) | 0.90 | 2.59% | (24m) | 0.93 | 2.25% | (2.0h) | |
| MOEA/D | 0.74 | 17.50% | (18h) | 0.66 | 30.28% | (12h) | 0.78 | 15.42% | (18h) | 0.69 | 27.63% | (12h) | |
| MBM | 0.86 | 5.21% | (13s) | 0.91 | 3.72% | (44s) | 0.90 | 2.46% | (13s) | 0.93 | 2.41% | (44s) | |
| MBM (aug) | 0.87 | 3.98% | (1.7m) | 0.91 | 3.28% | (5.9m) | 0.91 | 1.66% | (1.7m) | 0.93 | 2.10% | (5.9m) | |
| GMS-EB | 0.88 | 2.00% | (56s) | 0.93 | 1.81% | (9.1m) | 0.91 | 1.27% | (57s) | 0.94 | 1.33% | (9.2m) | |
| GMS-DH | 0.89 | 1.84% | (13s) | 0.92 | 2.11% | (49s) | 0.91 | 1.73% | (12s) | 0.93 | 2.13% | (52s) | |
| GMS-EB (aug) | 0.89 | 1.50% | (7.5m) | 0.93 | 1.47% | (1.2h) | 0.91 | 1.02% | (7.5m) | 0.94 | 1.09% | (1.2h) | |
| GMS-DH (aug) | 0.89 | 1.45% | (1.5m) | 0.93 | 1.61% | (6.6m) | 0.91 | 1.40% | (1.5m) | 0.93 | 1.92% | (6.6m) | |
Baselines: For the MOTSP and MGMOTSP, we utilize linear scalarization together with strong single-objective methods in the form of LKH3 (Helsgaun 2017) and Google OR-tools. We also benchmark against four Multi-Objective Evolutionary Algorithms (MOEAs): MOGLS (Ishibuchi and Murata 1998), MOEA/D (Zhang and Li 2007), NSGA-II (Deb et al. 2002) and NSGA-III (Blank, Deb, and Roy 2019). Below, we only show the results for the MOEA that performs best in each scenario. Finally, we design a neural baseline using the MatNet-architecture (Kwon et al. 2021) as encoder and a preference-conditioned MP network as decoder. We call this MatNet-Based Model (MBM) and our motivation is that this is a learning-based method that can be used almost ”off-the-shelf” by combining components from literature. In the multigraph setting, this architecture requires us to pre-process the input by pruning edges until we obtain a simple graph, which we do using linear scalarization.
In Appendix A and D, we provide further details about the neural benchmark and MOEAs respectively. In Appendix E, we also show results for all MOEAs as well as some other weaker heuristics.
Model Settings: We train all models for epochs using randomly generated instances per epoch with the ADAM optimizer (Kingma and Ba 2015) and a learning-rate of . For GMS-EB we let the number of layers be , while GMS-DH has and . For both models we use as the embedding dimension and attention heads in the GREAT layers and MP decoders. Regarding the number of samples in the decoding, we set for the selection-decoder during training and during inference. In the routing-decoder, we always set to the problem size as customary in the POMO framework (Kwon et al. 2020). Finally, we augment the models during inference with a factor of . For MBM, this is done using the original MatNet-augmentation (Kwon et al. 2021), while for GMS we use the scaling-augmentation of Lischka et al. (2024).
Metrics: We utilize the Hypervolume (HV) metric (Zitzler et al. 2003) to evaluate the performance of the methods. Formally, for a given set of points and a reference point that is dominated by all points in , the HV is defined as:
where denotes the Lebesgue measure in , and represents the axis-aligned hyperrectangle bounded by and . Intuitively, a higher HV generally indicates better performance.
In our experiments, we report the normalized HV, which is scaled to lie in the interval , averaged over 200 test instances. For the neural models we use 101 preferences during inference, which is also the number of subproblems for the benchmarks relying on scalarization. Apart from the HV, we provide the HV gap compared to the best-performing method as well as the inference time. In Appendix B, we also report some training times for the learning-based methods. For these methods, we used one NVIDIA Tesla A40 GPU with 48GB VRAM during training and inference. The non-learning-based methods were all tested using an Intel Xeon Gold 6338 CPU and 8 parallel processes.
5.2 Performance on Benchmark Problems
We show the performance of GMS together with the benchmarks in Table 1. As can be seen, GMS outperforms all benchmarks except LKH on the MOTSP and MGMOTSP, while it is the best performing model on the MOCVRP. This showcases that our architectures are generally applicable across a variety of distributions, instance sizes and problems. We note that the runtime for GMS-EB scales poorly with the problem size, but it is still fast compared to most benchmarks, especially without augmentation.
| FLEX2-20 | FLEX2-50 | |||||
| HV | Gap | Time | HV | Gap | Time | |
| MOEA/D | 0.75 | 9.19% | (2.7h) | 0.60 | 32.79% | (34h) |
| MBM | 0.72 | 12.89% | (4.9s) | 0.80 | 11.09% | (14s) |
| MBM (aug) | 0.73 | 11.98% | (34s) | 0.81 | 10.10% | (1.8m) |
| GMS-EB | 0.81 | 1.95% | (4.3s) | 0.89 | 0.97% | (60s) |
| GMS-DH | 0.80 | 3.37% | (3.1s) | 0.85 | 5.10% | (13s) |
| GMS-EB (aug) | 0.82 | 0.00% | (31s) | 0.90 | 0.00% | (7.9m) |
| GMS-DH (aug) | 0.82 | 0.33% | (16s) | 0.87 | 2.69% | (1.6m) |
5.3 Multigraph Problem without Clear Pruning
The main motivation behind learning to select edges in the multigraphs is to plan routes when no clear pruning method exists. Thus, we evaluate the performance on such a problem as well. Specifically, we consider an MGMOTSP with Time-Windows (MGMOTSPTW). Each edge is associated with a time and distance, while the objectives correspond to the number of violated time-windows and the total distance. Unlike the MGMOTSP with linear scalarization, it is not clear in this case which edges can be pruned apriori due to the more intricate time-window objective.
We display the performance on this problem in Table 2, and benchmark against MBM with LS pruning as well as MOEAs. Evidently, our models are the best performing ones, illustrating the value of learning to select edges in combination with learning-to-route.
5.4 Zero-Shot Generalization
We also test our models in a zero-shot setting on larger instances, which is displayed in Table 3. Both GMS-DH and GMS-EB remain competitive across both problems and all distributions.
| TMAT200 | XASY200 | ||||||
| HV | Gap | Time | HV | Gap | Time | ||
| MOTSP | LKH | 0.67 | 0.00% | (1.2h) | 0.95 | 0.00% | (1.3h) |
| OR Tools | 0.62 | 7.89% | (6.5h) | 0.90 | 4.46% | (4.0h) | |
| GMS-EB | 0.66 | 1.86% | (17m) | 0.92 | 2.78% | (17m) | |
| GMS-DH | 0.65 | 3.59% | (1.0m) | 0.90 | 5.22% | (1.0m) | |
| MGMOTSP | FLEX2-200 | FIX2-200 | |||||
| LKH | 0.97 | 0.00% | (1.2h) | 0.97 | 0.00% | (1.2h) | |
| OR Tools | 0.94 | 2.53% | (3.9h) | 0.95 | 1.73% | (4.0h) | |
| GMS-EB | 0.95 | 1.74% | (50m) | 0.96 | 1.34% | (50m) | |
| GMS-DH | 0.95 | 2.27% | (2.2m) | 0.94 | 2.56% | (2.2m) | |
5.5 Ablation Study
Finally, we showcase the ability of the selection-decoder in GMS-DH to select beneficial edges. Table 4 shows the result when replacing this component with a simple pruning function that removes edges with suboptimal linearly scalarized cost. For fairness, we also benchmark using both linearly scalarized reward and Chebyshev reward. As discussed, in the linear case the pruning should be optimal.
Evidently, Chebyshev reward improves the results substantially. Moreover, learning the selection compared to using a simple function also seems slightly better, even in the simple case of the MGMOTSP.
| FLEX2-50 | ||
| HV | Gap | |
| GMS-DH (Ch. Reward) | 0.8862 | 1.84% |
| GMS-DH (Lin. Reward) | 0.8781 | 2.74% |
| GMS-DH Simple (Lin. Reward) | 0.8789 | 2.65% |
| GMS-DH Simple (Ch. Reward) | 0.8847 | 2.00% |
6 Conclusion and Future Work
In this paper, we propose two methods to learn to construct solutions for routing problems on multigraphs. We highlight that while we solve problems with multiple objectives, our models can be valuable in all scenarios with no clear choice between alternative paths. In future works, we will apply our methods on more advanced problems.
A key weakness of our approach is that GMS-DH does not quite reach the same performance as GMS-EB in most cases. In this regard, there are works that are tangential to ours that could be utilized to obtain even better results for either GMS-DH or GMS-EB. Examples include the HV training of Lu et al. (2024) and diversity enhancement of Chen et al. (2023b). In the future, we will also explore decomposition methods like Ye et al. (2024) to extend our approach to even larger problem sizes.
Finally, from a broader perspective, we want to stress that routing problems on multigraphs are not well explored in literature. We believe one major reason is because of the complexity this formulation introduces. Learning-based methods can help alleviate this weakness and make the multigraph formulation more appealing.
Acknowledgements
This work was performed as a part of the research project “LEAR: Robust LEArning methods for electric vehicle Route selection” funded by the Swedish Electromobility Centre (SEC). The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Chalmers e-Commons partially funded by the Swedish Research Council through grant agreement no. 2022-06725. Through the affiliation of F.R., the work was also partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
Appendix A Architecture Details
Here, we highlight some of the main components in the model architectures that are not described in the main text.
A.1 Edge-Based Decoder in GMS-EB
In GMS-EB, we utilize an edge-based multi-pointer decoder which is based on Jin et al. (2023). Its mathematical formulation is the following. Given trainable weight matrices we form the query, representing the current routing context, as
Here, denotes an edge-embedding from the encoder, is the first visited edge and the most recently visited edge. Moreover, we utilize an enhanced context in which is the sum of all embeddings in the graph and is the sum over the visited edges according to
Let be the number of heads/pointers. The score for edge is formed according to
Note that the key is the embedding for the edge and is its dimension. Similarly to the original article, we subtract a cost to speed up the exploration and ensure that costly edges are considered less favorable from the start. In our case, the cost is multi-objective so we must scalarize it. We find that both Chebyshev and linear scalarization work well for this purpose. The last step is to form the logits as
Finally, softmax is applied to transform the logits to probabilities.
It must be noted that in our case, we preference-condition the decoder using a hyper-network. This is done by letting a network generate the weights and .
A.2 Selection-Decoder in GMS-DH
The task of this decoder is to prune parallel edges in the intermediate stage of GMS-DH, leaving only one between each node pair. Initially, we apply GREAT NB layers on the edge embeddings from the encoder to obtain embeddings specialized for this task. In the experiments we set . These layers are not preference conditioned. Hence, they do not need to be rerun for each preference. Next, given two nodes, we want to compute the probability of retaining each (directed) edge in between. For this purpose, we essentially utilize a multi-pointer score calculation similar to (6). Given a node pair , and the set of edges in between, , let the query be the average embedding according to
Moreover, let the key be the embedding . Then, form the score for edge as
where and . Again, note that we subtract a scalar cost to improve exploration. Also note that are generated by a hyper-network. Final edge probabilities are then computed using tanh-clipping and softmax as
A.3 MatNet-Based Benchmark Model
We visualize this architecture in Figure 3. The encoder consists of MatNet-layers (Kwon et al. 2021), while the decoder consists of an MP decoder (Jin et al. 2023) preference-conditioned with a hyper-network. The MatNet-architecture operates on bipartite graphs. It utilizes mixed-score attention layers together with fully-connected layers and normalization as in standard transformer layers (Vaswani et al. 2017). These attention layers ”mix” internally computed attention scores with external information from e.g. a distance matrix . To accommodate the use of several distance matrices, in our multi-objective setting, we follow the idea in Appendix A of the original paper. That is, we mix the internal attention score with all distance matrices using an MLP with inputs.
For the node-input to the initial MatNet-layer, in the absence of node-features we use zero-vectors for the ”row” nodes and one-hot vectors for the ”column” nodes, as in the original paper. If there are node features, e.g., demand in the MOCVRP, we use these as ”row” input while keeping the one-hot ”column” input. This allows us to augment each instance by sampling several one-hot combinations rather than scaling the instance as is done for GMS.
Lastly, we note that MatNet is not well-defined on multigraphs due to the fact that it utilizes the distance matrix to represent unique distances between nodes. It is possible to cast multigraphs into simple graphs, but not without increasing the number of nodes and edges drastically. Thus, we must pre-prune a multigraph before passing it through the encoder. This can be done with some sort of heuristic in the general case, but, as discussed, under LS and considering the MOTSP, a simple pruning method is to remove edges with sub-optimal linearly scalarized cost. Moreover, it can be noted that the encoder needs to be rerun for each new preference, which results in a slower architecture than if the encoding is done once.
Appendix B Model Training
Here we outline the curriculum learning which we briefly discussed in the main text. We also present an analysis of training performance.
B.1 Curriculum Learning Details
For both GMS-EB and GMS-DH we utilize a simple curriculum learning approach, with the purpose to decrease training times and to improve performance on large problems. The idea is simple: we train on small instances before training on large instances with the same distribution.
For instances with 20 nodes, we do not use any such curriculum. For instances with 50 nodes we start by training for 100 epochs on 20 nodes and then train on 50 nodes for 100 epochs. Finally, for instances with size 100, we start by training for 100 epochs on size 20. Then, for GMS-DH, we train for 50 epochs on size 50 and 50 epochs on size 100. As GMS-EB takes quite long to train for 100 nodes, we are satisfied with training for 100 epochs on size 50 for this model and then fine-tune it for 10 epochs on size 100, resulting in 210 epochs in total. Remark that our curriculum is quite arbitrary, but the model performance is not contingent on this exact setup.
Additionally, we start by training GMS-DH for epochs (included in the first 100 epochs) on the XASY distribution with 20 nodes. This is because the selection-decoder relies on a reward signal from the routing-decoder. As such, we find that performance and rate of convergence improves if the routing-decoder is pre-trained to obtain a stable and accurate signal. We also find that stability increases if we freeze the parameters of the shared encoder with respect to the loss of the selection-decoder. That is, only the loss from the routing-decoder is utilized to calculate encoder gradients.
B.2 Training Times
Table 5 shows the number of parameters for the models as well as training time per epoch for one problem on a Simple Graph (SG) and Multigraph (MG). The SG problem is the MOTSP with 50 nodes, while the MG problem is the MGMOTSP with 50 nodes and 2 parallel edges. It can be noted that GMS-EB takes a very long time to train, especially on multigraphs, while GMS-DH is more efficient and MBM is the most efficient. In fact, MBM is not affected by the multigraph representation as it works on pruned graphs.
We want to highlight that the training performance is a weakness of GMS. However, these models are very sample efficient, which is illustrated on examples with 20 nodes (that is, without curriculum learning) in Figure 4. Moreover, we try to remedy this weakness through the curriculum approach. When utilizing the curriculum learning, it is often enough to train on the small instance sizes to obtain a majority of performance and the rest is obtained after a few epochs of training on the larger instance. This can be seen in Figure 5, which showcases two examples with 50 nodes.
| #Params Enc. | #Params Dec. | Epoch SG | Epoch MG | |
| MBM | 2.38M | 266K | 4.7m | 4.7m |
| GMS-DH | 2.06M | 1.08M | 8.4m | 24m |
| GMS-EB | 1.98M | 266K | 51m | 1.7h |
Appendix C Routing Problems and Distributions
In this section, we detail the distributions and routing problems which we solve.
C.1 Distributions
The following distributions for node distances are used across the problems on simple graphs.
EUC - Euclidean. Here, we sample node coordinates uniformly and independently in the unit square. These are utilized to construct a distance matrix using the Euclidean distance. For a problem with node distances (e.g., the MOTSP with objectives) we sample pairs of coordinates independently.
TMAT - Asymmetric distance matrix obeying the triangle inequality (Kwon et al. 2021). The distance matrix is first populated with random independent elements and then post-processed until the triangle inequality is satisfied. For a problem with distances, we sample matrices independently.
XASY - Extremely asymmetric distance matrix (Gaile et al. 2022). The distance matrix does not satisfy the triangle inequality. It is obtained by independently sampling each element uniformly between 0 and 1. Again, we sample all matrices independently.
The following distributions are used for problems on multigraphs.
FLEX - Flexible number of edges between each node pair. Here we sample vectors of dimension representing edge attributes for edges between a node pair. Each attribute for each edge is sampled uniformly between 0 and 1. Then we remove dominated edges until only a Pareto set of edges remains.
FIX - Fixed number of edges between each node pair. We start by sampling vectors in the same way as the FLEX-distribution. Then we sort each attribute so that no edge is dominated. For , we sort one attribute in descending order and one attribute in ascending order.
Note that neither FLEX nor FIX obeys the triangle inequality, although such an assumption holds for real instances. We highlight that these distributions are utilized since they are simple and easy to sample from, creating a more transparent picture of the model performance. Motivated by the fact that XASY seems more difficult to learn than TMAT, we hypothesize that the triangle inequality would contribute positively to model performance.
C.2 MOTSP
Problem Description: In the multi-objective traveling salesman problem the goal is to visit all nodes/customers and return to the starting position. The Euclidean version of this problem is widely treated in the learning-for-routing community, see e.g., Lin, Yang, and Zhang (2022) and Fan et al. (2025) in the multi-objective case. In our case, we sample two distances, corresponding to the two objectives, between each pair of customers. The goal is to design a tour that minimizes the sum of the distances. A tour can start from any node.
Data Generation: Edge distances are sampled according to the EUC, TMAT and XASY distributions above.
Model Details: There are no node features for this problem, hence both the GREAT-based and MatNet-based encoders have the two distances as edge features, while the latter uses zero-vectors as ”row” input.
Evaluation Details: For all distributions, we utilize , and as reference for HV calculation for , and nodes respectively. In the zero-shot experiments, we have as reference for 150 nodes and for 200 nodes.
C.3 MOCVRP
Problem Description: In the multi-objective capacitated vehicle routing problem, a truck starts at a depot and must pick up goods from a specified number of customers. The vehicle has a fixed capacity while each customer has a demand. When the capacity is full, the vehicle must return to the depot. As many vehicles as required can be utilized to meet the demand of all customers.
Similarly to the TSP, this problem has been widely treated in the Euclidean setting using learning-based methods. As is common in the multi-objective setting (Lin, Yang, and Zhang 2022), we assume one objective is the traveled distance, while the other is the longest tour for a single vehicle (makespan). As such, each node-pair is only associated with one distance.
Data Generation: We assume that the capacity of the truck is 30, 40 and 50 for 20, 50 and 100 nodes respectively, while the demands are sampled uniformly from the set . This is the same setting as in e.g., Kool, van Hoof, and Welling (2018) and Lin, Yang, and Zhang (2022). Edge distances are sampled according to the EUC, TMAT and XASY distributions.
Model Details: In the decoding, we augment the routing context (represented by the query ) with the current load of the vehicle. For the GREAT-based encoders, we let the features of an edge be its distance as well as the demand of the customer at the end. On the other hand, the MatNet-based encoders only utilize the distance as an edge-feature while the demand is used as a node feature.
Evaluation Details: For all distributions, we utilize , and as reference for HV calculation for the three considered instance sizes.
C.4 MGMOTSP
Problem Description: As for the MOTSP, the goal of the multigraph MOTSP is to visit all nodes and return to the start. The only difference is that the underlying structure is a multigraph. Similarly, we define two distances for each node pair and these are associated with the objectives of the problem.
While this variant of the TSP on a multigraph has not been treated before to our knowledge (neither using learning nor other methods), some stochastic single-objective variants have been proposed in e.g., Tadei, Perboli, and Perfetti (2017) and Maggioni, Perboli, and Tadei (2014). Moreover, we believe that the MGMOTSP is a suitable benchmark problem for our solvers, as it is simple and easy to find good baselines for.
Data Generation: We sample distances between the nodes according to the FLEX and FIX distributions.
Model Details: As the MOTSP, the MGMOTSP lacks edge features. Hence, we utilize the same setup for the models as in the MOTSP. Moreover, for the MatNet-based model, we sparsify the underlying multigraph given each preference before it is fed into the encoder. This is done by removing all edges with sub-optimal LS cost compared to a parallel edge. To make this pruning strategy exact, we train MBM using linearly scalarized reward instead of the Chebyshev reward.
Evaluation Details: For the FLEX distribution, we utilize , and as reference for instance sizes 20, 50 and 100, regardless of . In the zero-shot experiments we set the references to and for FLEX with 150 and 200 nodes. For FIX, the reference is given by , where is the node count.
C.5 MGMOTSPTW
Problem Description: The goal in MGMOTSPTW is to create a tour that visits all customers and returns to the starting position. However, in this case we assume there is a fixed depot and that each customer/node is associated with a time-window. Preferably, the vehicle should arrive within the time-window for each customer. As such, we let one objective be the number of violated time-windows while the other is the traveled distance. Accordingly, each edge connecting two locations is associated with a travel time and a distance.
Multi-objective problems with similar objectives are fairly widespread in literature (Legrand et al. 2025), although not on multigraphs. In the multigraph framework, this problem is reminiscent of the vehicle routing problem with time-windows in Ben Ticha et al. (2017), but with a second objective that is the soft variant of the time-window constraint.
Data Generation: Distances and travel times are sampled according to the FLEX and FIX distributions. Additionally, we sample time-windows according to the ”medium” distribution of Chen et al. (2024) and Bi et al. (2024). In this setting, time-windows are sampled uniformly and independently, which results in difficult instances where there is a true trade-off between optimal paths and paths satisfying the time-window constraint. We also argue that this premise is simple for initial evaluation of our models, especially in comparison to e.g., the ”hard” instances of Chen et al. (2024), which tests out-of-distribution generalization.
Further, we assume that the vehicle can depart immediately from each customer, i.e., that there is no service time, and that no waiting occurs, even if a vehicle arrives before the starting time. The latter is for simplicity, as in general how long the vehicle should wait at a node is a decision in itself. We also argue that this assumption is fairly realistic.
Model Details: In the decoder, we enhance the routing context with an embedding of the current time. In the encoder, the node features in this case are the starting time and end time of the service, while the edge features are distance and travel time. Consequently, we augment the input features of GREAT so that an edge is associated with its distance, travel time and the time-window of the ending node. As usual for MatNet, the node features are used as ”row” inputs while the edge features are used in the distance matrix.
Note that in this case, pruning of the multigraph for MatNet is not obvious. For instance, removing an edge with a high travel time but low distance might not be optimal if traversing that edge would result in no violated time-windows anyway. Consequently, we apply the same simple pruning method as for the MGMOTSP. That is, we form the cost of each edge using a linear combination of its travel time and distance, then we remove all edges apart from the best one. We also choose the linear reward instead of the Chebyshev reward for MBM.
Evaluation Details: In this case, we let the reference value corresponding to the time-window violation be , while the reference value corresponding to the distance is the same as in the MGMOTSP case.
Appendix D Setup for MOEAS
As a base for the implementation of NSGA-II, NSGA-III and MOEA/D, we utilized Pymoo222https://pymoo.org/ and as a base for MOGLS, we utilized GeneticLocalSearch_TSP333https://github.com/kevin031060/Genetic˙Local˙Search˙TSP. For the former three, we limit the number of iterations to while the latter is limited to . Additionally, local search strategies is a necessity for efficient usage of genetic algorithms in combinatorial optimization. For the MOTSP, we utilized the 2-opt heuristic of Jaszkiewicz (2002), while we utilized the problem-specific local search and chromosome decoding strategy of Lacomme, Prins, and Sevaux (2006) for the MOCVRP.
However, in both cases we find that by accepting the best local move instead of just an improving one, the performance on large asymmetric problems improves substantially. As this affects the runtime negatively (due to the need to search through all moves), we limit the number of local moves in each search depending on the problem size. Additionally, for the algorithms not requiring preference weights for each individual, we must calculate such weights to compare the improvements. In the bi-objective case, this is done using the strategy of Lacomme, Prins, and Sevaux (2006) according to
where is the solution being evaluated and , are minimum and maximum values for the respective objectives.
In the multigraph setting, we design our own setup to tailor the MOEAS. To start with, inspired by one representation from Beke, Weiszer, and Chen (2021) regarding shortest path problems, we let a chromosome for a tour be given by a permutation of the nodes multiplied with added with the index of the edge to the next node. For example, in an MGMOTSP with 2 parallel edges and 5 nodes, a chromosome might be
For the mutation operator, we randomly either change one edge or reverse a segment of the tour. For the crossover, we utilize an edge recombination operator. Regarding the local search, we use a variant of 2-opt where we again accept the best local move. A local move is defined by a standard 2-opt move with the nodes, together with new edge selections given the changed node order. This edge selection is found by choosing the edge with the lowest LS cost between nodes that were previously not neighbors or have changed order.
We want to stress that the focus in this study is not to design an evolutionary algorithms, we simply want a baseline to compare against. We find that the proposed approach is somewhat efficient for small node counts, but performs quite badly for large graphs.
Appendix E Complementary Numerical Results
In this section, we show additional numerical results for more distributions, benchmarks and instance sizes.
E.1 Problems on Simple Graphs
Table 6 and 7 show results for the MOTSP and MOCVRP respectively. Apart from the benchmarks in the main text, we present the performance of three simple heuristics: Nearest Neighbor (NN), Nearest Insertion (NI) and Farthest Insertion (FI). As LKH and OR-tools, these utilize the linear scalarization to transform the MO problem into single-objective subproblems. Additionally, we show results on Euclidean instances (EUC). For these, there are many existing models in literature and thus we replace MBM with the Conditional Neural Heuristic (CNH) of Fan et al. (2025). This model is augmented with the Euclidean augmentation of Lin, Yang, and Zhang (2022).
On the EUC distribution, our models are generally slightly worse than CNH. This is to be expected though, as a common pattern in previous studies is that architectures built for directed graph inputs perform worse on EUC than those specialized for coordinate inputs (Lischka et al. 2024; Kwon et al. 2021).
| EUC20 | EUC50 | EUC100 | |||||||
| HV | Gap | Time | HV | Gap | Time | HV | Gap | Time | |
| LKH | 0.52 | 0.00% | (2.6m) | 0.58 | 0.00% | (14m) | 0.68 | 0.00% | (1.2h) |
| OR Tools | 0.51 | 0.97% | (2.0m) | 0.57 | 1.83% | (15m) | 0.67 | 1.57% | (1.3h) |
| NN | 0.46 | 11.44% | (1.9s) | 0.51 | 11.42% | (7.1s) | 0.63 | 8.23% | (25s) |
| NI | 0.47 | 8.53% | (11s) | 0.52 | 9.97% | (1.8m) | 0.63 | 7.82% | (12m) |
| FI | 0.50 | 3.44% | (13s) | 0.55 | 5.79% | (2.1m) | 0.65 | 4.96% | (14m) |
| MOGLS | 0.51 | 0.79% | (1.0h) | 0.58 | 0.74% | (3.8h) | 0.68 | 1.16% | (32h) |
| NSGA-II | 0.50 | 2.63% | (44m) | 0.52 | 11.18% | (3.4h) | 0.54 | 21.13% | (12h) |
| NSGA-III | 0.50 | 3.56% | (46m) | 0.50 | 14.07% | (3.4h) | 0.53 | 21.82% | (12h) |
| MOEA/D | 0.51 | 0.87% | (56h) | 0.57 | 2.53% | (3.6h) | 0.63 | 8.08% | (13h) |
| CNH | 0.51 | 1.16% | (6.5s) | 0.57 | 1.84% | (7.1s) | 0.67 | 1.89% | (25s) |
| CNH (aug) | 0.51 | 0.85% | (17s) | 0.57 | 1.26% | (44s) | 0.67 | 1.45% | (3.1m) |
| GMS-EB | 0.51 | 0.93% | (2.0s) | 0.57 | 1.98% | (25s) | 0.67 | 2.34% | (3.6m) |
| GMS-DH | 0.51 | 1.30% | (1.1s) | 0.57 | 2.64% | (3.8s) | 0.67 | 2.49% | (20s) |
| GMS-EB (aug) | 0.51 | 0.81% | (14s) | 0.57 | 1.57% | (3.3m) | 0.67 | 2.09% | (28m) |
| GMS-DH (aug) | 0.51 | 0.93% | (5.3s) | 0.57 | 1.91% | (27s) | 0.67 | 2.24% | (2.6m) |
| TMAT20 | TMAT50 | TMAT100 | |||||||
| LKH | 0.58 | 0.00% | (2.3m) | 0.58 | 0.00% | (6.4m) | 0.63 | 0.00% | (29m) |
| OR Tools | 0.56 | 3.01% | (3.3m) | 0.54 | 6.13% | (29m) | 0.59 | 7.12% | (2.5h) |
| NN | 0.48 | 16.66% | (1.9s) | 0.47 | 18.34% | (7.0s) | 0.54 | 15.25% | (25s) |
| NI | 0.53 | 8.78% | (11s) | 0.50 | 13.17% | (1.8m) | 0.55 | 13.24% | (12m) |
| FI | 0.54 | 6.60% | (13s) | 0.52 | 10.65% | (2.0m) | 0.57 | 10.84% | (14m) |
| MOGLS | 0.54 | 7.72% | (20m) | 0.40 | 30.90% | (2.7h) | 0.38 | 40.03% | (19h) |
| NSGA-II | 0.57 | 2.17% | (44m) | 0.51 | 12.91% | (3.5h) | 0.47 | 25.98% | (12h) |
| NSGA-III | 0.57 | 2.87% | (45m) | 0.50 | 14.54% | (3.5h) | 0.46 | 27.53% | (13h) |
| MOEA/D | 0.57 | 1.43% | (56m) | 0.53 | 9.51% | (3.7h) | 0.50 | 20.93% | (13h) |
| MBM | 0.57 | 2.29% | (1.3s) | 0.50 | 13.51% | (3.9s) | 0.56 | 11.16% | (19s) |
| MBM (aug) | 0.58 | 0.96% | (5.6s) | 0.52 | 10.17% | (27s) | 0.58 | 9.17% | (2.4m) |
| GMS-EB | 0.58 | 0.88% | (2.2s) | 0.57 | 1.50% | (25s) | 0.63 | 1.45% | (3.6m) |
| GMS-DH | 0.58 | 1.08% | (1.3s) | 0.57 | 2.36% | (4.1s) | 0.62 | 2.76% | (20s) |
| GMS-EB (aug) | 0.58 | 0.67% | (15s) | 0.57 | 1.14% | (3.3m) | 0.63 | 1.13% | (29m) |
| GMS-DH (aug) | 0.58 | 0.79% | (5.6s) | 0.57 | 1.76% | (28s) | 0.62 | 2.19% | (2.6m) |
| XASY20 | XASY50 | XASY100 | |||||||
| LKH | 0.75 | 0.00% | (2.3m) | 0.83 | 0.00% | (6.9m) | 0.90 | 0.00% | (32m) |
| OR Tools | 0.71 | 5.49% | (3.0m) | 0.77 | 7.25% | (24m) | 0.85 | 6.06% | (1.8h) |
| NN | 0.60 | 20.11% | (1.8s) | 0.70 | 16.32% | (7.0s) | 0.80 | 10.94% | (25s) |
| NI | 0.61 | 17.76% | (11s) | 0.66 | 21.37% | (1.8m) | 0.73 | 18.88% | (12m) |
| FI | 0.60 | 19.57% | (12s) | 0.64 | 23.62% | (2.1m) | 0.72 | 20.42% | (14m) |
| MOGLS | 0.65 | 13.23% | (19m) | 0.51 | 39.16% | (2.6h) | 0.47 | 47.82% | (19h) |
| NSGA-II | 0.72 | 4.06% | (44m) | 0.70 | 15.63% | (3.5h) | 0.64 | 29.74% | (13h) |
| NSGA-III | 0.71 | 4.56% | (45m) | 0.70 | 16.64% | (3.5h) | 0.63 | 30.13% | (13h) |
| MOEA/D | 0.73 | 2.68% | (56m) | 0.73 | 12.68% | (3.6h) | 0.70 | 22.22% | (13h) |
| MBM | 0.68 | 9.10% | (1.4s) | 0.75 | 10.37% | (3.7s) | 0.84 | 7.41% | (19s) |
| MBM (aug) | 0.70 | 5.86% | (5.5s) | 0.76 | 8.28% | (27s) | 0.85 | 6.34% | (2.4m) |
| GMS-EB | 0.73 | 1.95% | (2.3s) | 0.81 | 2.94% | (25s) | 0.88 | 3.03% | (3.6m) |
| GMS-DH | 0.73 | 2.37% | (1.9s) | 0.80 | 4.12% | (3.9s) | 0.86 | 4.87% | (20s) |
| GMS-EB (aug) | 0.74 | 1.30% | (15s) | 0.82 | 2.06% | (3.3m) | 0.88 | 2.76% | (28m) |
| GMS-DH (aug) | 0.74 | 1.57% | (5.8s) | 0.81 | 2.84% | (28s) | 0.87 | 3.85% | (2.6m) |
| EUC20 | EUC50 | EUC100 | |||||||
| HV | Gap | Time | HV | Gap | Time | HV | Gap | Time | |
| MOGLS | 0.23 | 3.35% | (4.1h) | 0.17 | 37.39% | (26h) | 0.00 | 100.00% | (108h) |
| NSGA-II | 0.23 | 0.13% | (2.6h) | 0.27 | 2.14% | (24h) | 0.20 | 18.91% | (75h) |
| NSGA-III | 0.23 | 0.21% | (2.7h) | 0.26 | 2.91% | (25h) | 0.20 | 21.23% | (75h) |
| MOEA/D | 0.23 | 0.90% | (2.9h) | 0.26 | 2.73% | (25h) | 0.23 | 7.40% | (73h) |
| CNH | 0.23 | 0.77% | (5.0s) | 0.27 | 0.59% | (9.5s) | 0.25 | 0.56% | (29s) |
| CNH (aug) | 0.23 | 0.00% | (29s) | 0.27 | 0.00% | (58s) | 0.25 | 0.00% | (3.6m) |
| GMS-EB | 0.23 | 0.30% | (3.4s) | 0.27 | 0.44% | (33s) | 0.25 | 0.56% | (4.1m) |
| GMS-DH | 0.23 | 1.33% | (2.3s) | 0.27 | 1.11% | (6.6s) | 0.25 | 1.36% | (30s) |
| GMS-EB (aug) | 0.23 | 0.13% | (22s) | 0.27 | 0.18% | (4.7m) | 0.25 | 0.32% | (54m) |
| GMS-DH (aug) | 0.23 | 0.90% | (10s) | 0.27 | 0.66% | (46s) | 0.25 | 0.92% | (4.1m) |
| TMAT20 | TMAT50 | TMAT100 | |||||||
| MOGLS | 0.40 | 1.78% | (4.1h) | 0.31 | 35.06% | (28h) | 0.05 | 89.38% | (116h) |
| NSGA-II | 0.40 | 1.38% | (2.7h) | 0.36 | 23.04% | (23h) | 0.13 | 71.72% | (76h) |
| NSGA-III | 0.40 | 1.98% | (2.7h) | 0.35 | 25.91% | (24h) | 0.12 | 74.11% | (76h) |
| MOEA/D | 0.39 | 2.59% | (3.0h) | 0.41 | 13.98% | (24h) | 0.19 | 58.70% | (76h) |
| MBM | 0.40 | 1.14% | (2.4s) | 0.47 | 1.33% | (6.2s) | 0.43 | 4.50% | (27s) |
| MBM (aug) | 0.40 | 0.86% | (10s) | 0.47 | 0.99% | (41s) | 0.44 | 3.29% | (3.3m) |
| GMS-EB | 0.40 | 0.25% | (3.5s) | 0.47 | 0.25% | (36s) | 0.45 | 0.57% | (4.2m) |
| GMS-DH | 0.40 | 0.64% | (2.5s) | 0.47 | 0.76% | (6.8s) | 0.45 | 1.46% | (32s) |
| GMS-EB (aug) | 0.41 | 0.00% | (23s) | 0.47 | 0.00% | (5.2m) | 0.45 | 0.00% | (33m) |
| GMS-DH (aug) | 0.40 | 0.32% | (10s) | 0.47 | 0.36% | (47s) | 0.45 | 0.79% | (4.3m) |
| XASY20 | XASY50 | XASY100 | |||||||
| MOGLS | 0.59 | 2.39% | (4.0h) | 0.49 | 34.85% | (29h) | 0.13 | 83.50% | (121h) |
| NSGA-II | 0.60 | 1.24% | (2.7h) | 0.55 | 27.19% | (24h) | 0.28 | 64.39% | (85h) |
| NSGA-III | 0.59 | 2.21% | (2.7h) | 0.52 | 31.06% | (25h) | 0.25 | 68.40% | (86h) |
| MOEA/D | 0.59 | 3.24% | (3.0h) | 0.65 | 13.88% | (24h) | 0.42 | 47.23% | (83h) |
| MBM | 0.57 | 5.95% | (2.3s) | 0.70 | 6.57% | (6.2s) | 0.60 | 24.47% | (27s) |
| MBM (aug) | 0.59 | 3.06% | (10s) | 0.72 | 4.13% | (42s) | 0.64 | 19.60% | (3.4m) |
| GMS-EB | 0.60 | 1.44% | (3.5s) | 0.74 | 1.89% | (36s) | 0.79 | 1.23% | (4.3m) |
| GMS-DH | 0.58 | 3.40% | (2.4s) | 0.72 | 4.13% | (6.6s) | 0.77 | 3.60% | (33s) |
| GMS-EB (aug) | 0.61 | 0.00% | (23s) | 0.75 | 0.00% | (5.0m) | 0.80 | 0.00% | (48m) |
| GMS-DH (aug) | 0.60 | 1.29% | (10s) | 0.74 | 1.65% | (47s) | 0.79 | 1.21% | (4.7m) |
E.2 Multigraph Problems
We display additional results for the MGMOTSP and MGMOTSPTW in Tables 8-10. As for the MOTSP, we include NN, NI and FI for MGMOTSP. In the case of MGMOTSPTW, we also include results for LKH and OR-tools when treating the problem as the MGMOTSP. That is, we neglect the time-windows and treat the time duration of each edge exactly as the distance. We call these methods implicit LKH and implicit OR-tools, as they implicitly minimize the time-window violations by reducing the time the tour takes.
| FLEX2-20 | FLEX2-50 | FLEX2-100 | |||||||
| HV | Gap | Time | HV | Gap | Time | HV | Gap | Time | |
| LKH | 0.85 | 0.00% | (1.2m) | 0.90 | 0.00% | (6.3m) | 0.94 | 0.00% | (31m) |
| OR Tools | 0.82 | 3.30% | (3.0m) | 0.86 | 4.21% | (24m) | 0.91 | 3.46% | (1.8h) |
| NN | 0.74 | 13.34% | (1.9s) | 0.81 | 10.38% | (7.1s) | 0.88 | 6.71% | (25s) |
| NI | 0.75 | 11.13% | (9.6s) | 0.78 | 13.13% | (1.7m) | 0.84 | 11.32% | (12m) |
| FI | 0.75 | 11.98% | (12s) | 0.77 | 14.30% | (2.0m) | 0.83 | 12.20% | (14m) |
| MOGLS | 0.71 | 16.29% | (1.9h) | 0.49 | 45.84% | (29h) | 0.43 | 54.91% | (18h) |
| NSGA-II | 0.81 | 5.06% | (1.3h) | 0.74 | 17.62% | (19h) | 0.64 | 32.11% | (15h) |
| NSGA-III | 0.80 | 5.82% | (1.3h) | 0.74 | 18.48% | (19h) | 0.63 | 33.16% | (14h) |
| MOEA/D | 0.80 | 6.02% | (1.4h) | 0.74 | 17.50% | (18h) | 0.66 | 30.28% | (12h) |
| MBM | 0.82 | 3.30% | (4.4s) | 0.86 | 5.21% | (13s) | 0.91 | 3.72% | (44s) |
| MBM (aug) | 0.84 | 1.50% | (32s) | 0.87 | 3.98% | (1.7m) | 0.91 | 3.28% | (5.9m) |
| GMS-EB | 0.83 | 1.61% | (4.0s) | 0.88 | 2.00% | (56s) | 0.93 | 1.81% | (9.1m) |
| GMS-DH | 0.84 | 1.34% | (3.2s) | 0.89 | 1.84% | (13s) | 0.92 | 2.11% | (49s) |
| GMS-EB (aug) | 0.84 | 0.95% | (34s) | 0.89 | 1.50% | (7.5m) | 0.93 | 1.47% | (1.2h) |
| GMS-DH (aug) | 0.84 | 0.86% | (19s) | 0.89 | 1.45% | (1.5m) | 0.93 | 1.61% | (6.6m) |
| FIX2-20 | FIX2-50 | FIX2-100 | |||||||
| HV | Gap | Time | HV | Gap | Time | HV | Gap | Time | |
| LKH | 0.85 | 0.00% | (1.2m) | 0.92 | 0.00% | (6.3m) | 0.95 | 0.00% | (30m) |
| OR Tools | 0.83 | 2.42% | (2.9m) | 0.90 | 2.59% | (24m) | 0.93 | 2.25% | (2.0h) |
| NN | 0.77 | 9.86% | (1.9s) | 0.86 | 6.34% | (7.1s) | 0.91 | 4.23% | (25s) |
| NI | 0.78 | 8.39% | (9.6s) | 0.85 | 7.95% | (1.7m) | 0.88 | 7.14% | (12m) |
| FI | 0.78 | 9.23% | (11s) | 0.84 | 8.82% | (2.0m) | 0.88 | 7.89% | (14m) |
| MOGLS | 0.74 | 13.68% | (1.9h) | 0.65 | 29.64% | (29h) | 0.52 | 45.00% | (19h) |
| NSGA-II | 0.81 | 5.64% | (1.3h) | 0.77 | 16.56% | (19h) | 0.69 | 27.94% | (15h) |
| NSGA-III | 0.80 | 6.85% | (1.3h) | 0.75 | 18.79% | (19h) | 0.66 | 30.49% | (15h) |
| MOEA/D | 0.80 | 6.20% | (1.4h) | 0.78 | 15.42% | (18h) | 0.69 | 27.63% | (12h) |
| MBM | 0.84 | 1.57% | (4.4s) | 0.90 | 2.46% | (13s) | 0.93 | 2.41% | (44s) |
| MBM (aug) | 0.85 | 0.50% | (32s) | 0.91 | 1.66% | (1.7m) | 0.93 | 2.10% | (5.9m) |
| GMS-EB | 0.84 | 1.46% | (3.9s) | 0.91 | 1.27% | (57s) | 0.94 | 1.33% | (9.2m) |
| GMS-DH | 0.84 | 1.37% | (3.2s) | 0.91 | 1.73% | (12s) | 0.93 | 2.13% | (52s) |
| GMS-EB (aug) | 0.85 | 0.99% | (28s) | 0.91 | 1.02% | (7.5m) | 0.94 | 1.09% | (1.2h) |
| GMS-DH (aug) | 0.85 | 0.82% | (18s) | 0.91 | 1.40% | (1.5m) | 0.93 | 1.92% | (6.6m) |
| FLEX2-20 | FLEX2-50 | |||||
| HV | Gap | Time | HV | Gap | Time | |
| Implicit LKH | 0.27 | 66.74% | (1.1m) | 0.14 | 84.11% | (6.7m) |
| Implicit OR Tools | 0.24 | 70.78% | (3.4m) | 0.13 | 85.34% | (25m) |
| MOGLS | 0.64 | 22.92% | (2.7h) | 0.35 | 60.96% | (21h) |
| NSGA-II | 0.69 | 16.55% | (2.4h) | 0.51 | 43.47% | (35h) |
| NSGA-III | 0.69 | 16.01% | (2.5h) | 0.50 | 44.53% | (36h) |
| MOEA/D | 0.75 | 9.19% | (2.7h) | 0.60 | 32.79% | (34h) |
| MBM | 0.72 | 12.89% | (4.9s) | 0.80 | 11.09% | (14s) |
| MBM (aug) | 0.73 | 11.98% | (34s) | 0.81 | 10.10% | (1.8m) |
| GMS-EB | 0.81 | 1.95% | (4.3s) | 0.89 | 0.97% | (60s) |
| GMS-DH | 0.80 | 3.37% | (3.1s) | 0.85 | 5.10% | (13s) |
| GMS-EB (aug) | 0.82 | 0.00% | (31s) | 0.90 | 0.00% | (7.9m) |
| GMS-DH (aug) | 0.82 | 0.33% | (16s) | 0.87 | 2.69% | (1.6m) |
E.3 Generalization to Larger Instances
In Tables 11-12, we report the generalization performance zero-shot for our methods in the MOTSP and MGMOTSP cases. It can be seen that both models retain competitive performance. In particular, GMS-EB often performs close to LKH while being faster.
| EUC150 | EUC200 | |||||
| HV | Gap | Time | HV | Gap | Time | |
| LKH | 0.73 | 0.00% | (1.6h) | 0.76 | 0.00% | (3.0h) |
| OR Tools | 0.72 | 1.38% | (1.7h) | 0.75 | 1.24% | (3.5h) |
| CNH ZS | 0.72 | 2.04% | (50s) | 0.74 | 2.22% | (1.3m) |
| GMS-EB ZS | 0.71 | 2.71% | (6.7m) | 0.74 | 2.93% | (17m) |
| GMS-DH ZS | 0.71 | 2.82% | (30s) | 0.74 | 3.12% | (1.0m) |
| TMAT150 | TMAT200 | |||||
| LKH | 0.65 | 0.00% | (39m) | 0.67 | 0.00% | (1.2h) |
| OR Tools | 0.60 | 7.57% | (3.3h) | 0.62 | 7.89% | (6.5h) |
| GMS-EB ZS | 0.64 | 1.64% | (6.7m) | 0.66 | 1.86% | (17m) |
| GMS-DH ZS | 0.63 | 3.20% | (30s) | 0.65 | 3.59% | (1.0m) |
| XASY150 | XASY200 | |||||
| LKH | 0.93 | 0.00% | (40m) | 0.95 | 0.00% | (1.3h) |
| OR Tools | 0.88 | 5.12% | (2.2h) | 0.90 | 4.46% | (4.0h) |
| GMS-EB ZS | 0.90 | 2.77% | (6.7m) | 0.92 | 2.78% | (17m) |
| GMS-DH ZS | 0.88 | 5.10% | (30s) | 0.90 | 5.22% | (1.0m) |
| FLEX2-150 | FLEX2-200 | |||||
| HV | Gap | Time | HV | Gap | Time | |
| LKH | 0.96 | 0.00% | (39m) | 0.97 | 0.00% | (1.2h) |
| OR Tools | 0.93 | 2.93% | (2.1h) | 0.94 | 2.53% | (3.9h) |
| GMS-EB ZS | 0.94 | 1.78% | (18m) | 0.95 | 1.74% | (50m) |
| GMS-DH ZS | 0.94 | 2.17% | (1.0m) | 0.95 | 2.27% | (2.2m) |
| FIX2-150 | FIX2-200 | |||||
| LKH | 0.96 | 0.00% | (39m) | 0.97 | 0.00% | (1.2h) |
| OR Tools | 0.94 | 1.96% | (2.2h) | 0.95 | 1.73% | (4.0h) |
| GMS-EB ZS | 0.95 | 1.35% | (18m) | 0.96 | 1.34% | (50m) |
| GMS-DH ZS | 0.94 | 2.38% | (1.1m) | 0.94 | 2.56% | (2.2m) |

### v2

@@
-Learning to Solve Multi-Objective Routing Problems on Multigraphs
+Neural Approaches for Multi-Objective Routing on Multigraphs
 Abstract
 Learning-based methods for routing have gained significant attention in recent years, both in single-objective and multi-objective contexts.
-However, the multigraph setting, where multiple paths with distinct attributes can exist between destinations, has largely been overlooked, despite its high practical relevancy.
-In this paper, we introduce two neural approaches to address multi-objective routing on multigraphs.
-Our first approach works directly on the multigraph, by autoregressively selecting edges until a tour is completed.
-On the other hand, our second model first prunes the multigraph into a simple graph and then builds routes.
-We validate both models experimentally and find that they demonstrate strong performance across a variety of problems, including the Traveling Salesman Problem (TSP) and Capacitated Vehicle Routing Problem (CVRP)111Our code will be made publicly available upon acceptance of the paper..
+Yet, existing methods are unsuitable for routing on multigraphs, which feature multiple edges with distinct attributes between node pairs, despite their strong relevance in real-world scenarios.
+In this paper, we propose two graph neural network-based methods to address multi-objective routing on multigraphs.
+Our first approach operates directly on the multigraph by autoregressively selecting edges until a tour is completed.
+The second model first simplifies the multigraph via a learned pruning strategy and then performs routing on the resulting simple graph.
+We evaluate both models empirically and demonstrate their strong performance across a range of problems and distributions111Our code will be released publicly upon paper acceptance..
 1 Introduction
-Routing problems, including the well-known Traveling Salesman Problem (TSP), are central for optimizing operations in logistics, delivery and beyond.
-However, these problems are generally NP-hard and difficult to solve exactly in practical, real-time settings.
-While exact solvers provide high-quality solutions, they do not scale well.
-At the same time, hand-crafted heuristics are often efficient but time-consuming to design and typically lack generalizability across problem variants.
-Recently, neural solvers have emerged as a promising alternative, offering speed, accuracy, and broad applicability across routing problems (Kool, van Hoof, and Welling 2018; Ye et al.
-2024; Drakulic, Michel, and Andreoli 2025).
-However, most learning-based approaches focus on Euclidean settings, where customers are points in a plane and travel costs are based on the Euclidean distance.
-Real-world road networks are far more complex and travel cost is usually asymmetric, affected by elevation, one-way streets and traffic.
-Further, when there are multiple conflicting criteria of importance, there can be multiple paths between customers that cannot be chosen between apriori.
-In such cases, a directed multigraph, allowing multiple, distinct paths between the same locations, is a more realistic and powerful input-structure.
-In fact, using this representation has empirically been shown to increase solution quality significantly compared to the simple graph representation (Ben Ticha et al.
-2017).
-Yet this comes at the cost of increased complexity, underscoring the need for new methods capable of handling multigraph-based routing efficiently.
-The challenge intensifies in Multi-Objective (MO) routing problems, such as the Multi-Objective TSP (MOTSP), where several criteria (e.g., distance and time) must be optimized simultaneously.
-Even without multigraphs, MO problems are hard: not only because their single-objective counter-part is often NP-hard, but also because the number of Pareto optimal solutions can increase exponentially with the size of the problem (Ehrgott and Gandibleux 2000).
-Pareto Set Learning (PSL) has recently emerged as a powerful framework for MO routing.
-By training offline on similar instances, PSL approximates the Pareto set quickly during inference.
-Moreover, since the framework is based on reinforcement learning, there is no need for oracle solvers for supervised training.
-These methods have shown strong results in Euclidean settings (Lin, Yang, and Zhang 2022; Fan et al.
-2025), but there is currently no general PSL framework for routing on multigraphs or in non-Euclidean settings, despite their higher practical relevance.
-We attempt to bridge this gap. Concretely, our main contributions are:
-We extend PSL to handle multigraph inputs, where the model selects not only the next node but also the path taken.
-To our knowledge, this is the first neural solver designed for such input structures.
-Although we treat multi-objective optimization, we hypothesize that our approach can be useful in single-objective constrained and stochastic cases too.
-As a byproduct, our model constitutes the first general-purpose, multi-objective solver capable of handling a variety of routing problems on asymmetric graphs.
-In order to do this, we propose two separate models: One edge-based that is simple and accurate but relatively slow and one dual-task model that first sparsifies the multigraph into a simple graph and then selects routes.
-The latter is faster but more complex.
-Experimentally, we show that our approach achieves competitive results for several MO routing problems on asymmetric graphs and multigraphs, including the MOTSP, Multi-Objective Capacitated Vehicle Routing Problem (MOCVRP) and MOTSP with time-windows.
+Routing problems, including the well-known Traveling Salesman Problem (TSP), are central for optimizing operations in modern transport systems.
+A key assumption is often that an optimal path can be selected beforehand between each pair of locations.
+However, in reality, several paths may exist that cannot be chosen between a priori.
+This naturally leads to a multigraph representation.
+Notably, by planning on a multigraph rather than a simple graph, Ben Ticha et al.
+(2017) demonstrated a cost reduction of up to for the vehicle routing problem with time-windows.
+Similarly Lai, Caliskan Demirag, and Leung (2016) demonstrated average savings of circa for a heterogeneous vehicle routing problem.
+Such results can be pivotal in for instance last-mile-delivery, which accounts for of total supply chain cost according to Capgemini Research Institute (2023).
+Yet, routing problems on simple graphs are already difficult to solve, especially in real-time, as they are generally NP-hard.
+The multigraph setup further increases this complexity.
+Moreover, such a framework is particularly interesting in the Multi-Objective (MO) case, where each edge may be associated with several conflicting cost criteria, such as time, distance or energy.
+MO routing is typically harder, as the number of Pareto optimal solutions can grow exponentially with the size of the problem (Ehrgott and Gandibleux 2000).
+Clearly, MO routing on multigraphs poses significant challenges, while remaining highly relevant for real-world applications.
+Recently, however, learning-based methods have emerged as promising options for routing, offering speed, accuracy, and broad applicability across problems (Kool, van Hoof, and Welling 2018; Drakulic, Michel, and Andreoli 2025).
+Yet, current approaches are designed for simple graphs and not suited for the multigraph representation.
+This is especially true in the MO setting, where neural approaches have demonstrated excellent results on Euclidean instances (Lin, Yang, and Zhang 2022; Fan et al.
+2025), but lack support even for asymmetric graphs, in spite of their higher practical relevance (Boyacı, Dang, and Letchford 2021).
+We attempt to bridge the gap between learning-based methods for MO routing and accurate network representations. Concretely, our main contributions are:
+We design models to handle multigraph inputs, where not only the next node but also the edge must be selected.
+To our knowledge, we present the first neural solvers designed for such input structures.
+Although we deal with MO optimization, our approach can be highly useful in single-objective cases too.
+Our models are also capable of handling a variety of MO routing problems on asymmetric graphs.
+In contrast, current neural MO methods are either problem-specific or designed exclusively for the Euclidean setting, where edge costs are symmetric and determined only by the node coordinates.
+We propose two separate models based on Graph Neural Networks (GNNs): One edge-based working directly on the multigraph and one dual-task model that first prunes the multigraph into a simple graph and then selects routes.
+The former is simple but slow while the latter is more complex but faster.
+Experimentally, we show that our approach achieves competitive results for several MO routing problems on asymmetric graphs and multigraphs.
 2 Related Work
 2.1 Learning for Routing
-Motivated by the success of end-to-end learning rather than hand-crafted heuristics in fields such as image classification and language modeling, neural heuristics for routing have gained increasing attention during the last years.
-Broadly, methods can be classified by how they obtain the tours.
+Motivated by the success of end-to-end learning in fields such as image classification and language modeling, neural solutions for routing have gained increasing attention during the last years.
+Broadly, methods can be classified by how they obtain tours.
 Two prominent alternatives are autoregressive construction, based on iterative building of tours one node at a time, and non-autoregressive construction, that outputs heatmaps in a one-shot fashion for down-stream decoding with some search method.
-Early works in the first category include Vinyals, Fortunato, and Jaitly (2015), which utilized an oracle TSP solver for supervised learning, and Bello et al.
-(2017), which introduced reinforcement learning for training.
-By using attention models and inherent symmetries in the problem, first Kool, van Hoof, and Welling (2018) and then Kwon et al.
-(2020) obtained close to 0 optimality gap on various routing problems.
-Regarding asymmetric problems, Kwon et al.
-(2021) proposed MatNet, a mixed-attention architecture working on bi-partite graphs.
-More recent works are Jin et al.
-(2023); Drakulic et al.
-(2023); Chen et al.
-(2024); Bi et al.
+Examples from the first category include Vinyals, Fortunato, and Jaitly (2015); Kool, van Hoof, and Welling (2018); Kwon et al.
 (2020).
-These papers have improved results, can handle constraints better and can scale to even larger problems than previously.
-Another notable example is Drakulic, Michel, and Andreoli (2025), where one model learns to solve a multitude of combinatorial optimization problems, including examples from routing.
-In the second category, some notable works are Joshi, Laurent, and Bresson (2019); Fu, Qiu, and Zha (2021); Qiu, Sun, and Yang (2022); Sun and Yang (2023).
-Typically, these methods perform worse than autoregressive approaches, but scale better to large instances due to lower space and time complexity.
-An interesting recent work combining the two approaches to solve large instances (up to nodes for the TSP) is Ye et al.
-(2024).
-2.2 Multi-Objective Learning for Routing
-The literature on multi-objective routing is more sparse than in the single-objective case.
-Some early works (Li, Zhang, and Wang 2021; Wu, Wang, and Zhang 2020) utilize multiple models, where one model is responsible for one preference between the objectives.
-This quickly becomes infeasible as the number of models required grows exponentially with the number of objectives (Ehrgott 2005).
-Pareto-set learning for multi-objective routing using a single model was proposed in Lin, Yang, and Zhang (2022).
-They used the hyper-networks of Navon et al.
-(2021) to condition the decoder on the preference.
-Recent papers that feature improved results with a similar conditioning method include Fan et al.
-(2025); Lu et al.
-(2024).
-An alternative approach is to use meta-learning to train a single model that can be quickly fine-tuned over a few steps to solve individual subproblems.
-Two representative works in this direction are Zhang et al.
-(2021); Chen et al.
-(2023a).
-Regarding asymmetric problems in the context of multi-objective optimization, the only existing learning-based methods we are aware of are those proposed in Santiyuda et al.
+Examples from the second category are Joshi, Laurent, and Bresson (2019); Fu, Qiu, and Zha (2021); Sun and Yang (2023).
+For a more comprehensive discussion of these methods, we refer the reader to Appendix A, which presents additional related work.
+The literature on MO routing is more sparse than in the single-objective case.
+A key contribution was made by Lin, Yang, and Zhang (2022), who proposed an autoregressive construction approach using a single neural model to learn the Pareto set of solutions.
+Regarding asymmetric problems, the only existing learning-based methods we are aware of are those introduced by Santiyuda et al.
 (2024) and Zhou, Kong, and Wang (2025).
-Notably, both of these methods are tailored to particular problem types and evaluated within narrowly defined domains.
-In contrast, our approach is designed with a very general framework, capable of handling a broad class of routing problems.
+However, both these approaches are tailored to specific problems and have been evaluated within narrowly defined domains.
+In contrast, our method is capable of addressing a broad class of routing problems and is evaluated across various problems and settings.
 2.2 Routing on Multigraphs
-In the multi-objective case, a multigraph is often a more suitable representation of a road network than a simple graph.
-In fact, this network model has become more and more common for MO shortest path problems.
-One solution is genetic algorithms, see for instance Beke, Weiszer, and Chen (2021) for a comprehensive benchmarking study.
-Other studies have proposed A∗-related algorithms, e.g., AMOA∗ (Weiszer, Burke, and Chen 2020) and MOMGA∗ (Liu, Chen, and Weiszer 2022), to efficiently obtain a Pareto set of shortest paths.
-The multigraph framework has also been applied to routing problems, but to a lower extent.
-The first study in this regard was Garaix et al.
-(2010), who designed a simple heuristic and branch-and-price method for a single-objective on-demand transportation problem.
-Other works proposing either exact solvers or heuristics for single-objective constrained problems or multi-objective problems are Lai, Caliskan Demirag, and Leung (2016); Ben Ticha et al.
+The multigraph representation has become more and more common for MO shortest path problems during recent years (Weiszer, Burke, and Chen 2020; Beke, Weiszer, and Chen 2021).
+Such a framework has also been applied to combinatorial routing problems like Vehicle Routing Problems (VRPs), but to a lower extent.
+Garaix et al.
+(2010) performed the first study in this regard and later works include Lai, Caliskan Demirag, and Leung (2016); Ben Ticha et al.
 (2017); Tikani, Setak, and Demir (2021).
-In addition, a particularly interesting study is Ben Ticha et al.
-(2017), where it is clearly concluded that a multigraph setup yields significantly lower costs compared to a simple graph representation for the capacitated vehicle routing problem with time-windows.
-On real-world instances, a solution cost reduction of up to is reported.
-At the same time, the computation time increases significantly under this perspective, and the authors highlight the need for new heuristics.
-We believe learning-based methods can fill this niche, as they have proven to be efficient, effective and broadly applicable.
-To the best of our knowledge, this work presents the first learning-based routing approach to address problems on multigraphs.
-3 Multi-Objective Optimization
-A general multi-objective optimization problem can be written as
-where and .
-Unlike the Single-Objective (SO) case, the solution to such a problem is not one objective value with corresponding optimal solutions but a set of values known as the Pareto Front (PF) with corresponding solutions in the Pareto Set (PS).
-Formally, we define the PS as the set
+A common theme across these studies is that the multigraph representation leads to considerably improved solutions, while simultaneously increasing the complexity of these already difficult problems.
+Consequently, learning-based methods should be able to fill a clear need for new solvers in this area, as they have proven to be efficient, effective and broadly applicable.
+To the best of our knowledge, this work presents the first learning-based approach to address combinatorial routing problems such as the TSP and VRPs on multigraphs.
+3 Problem Formulation
+We define a general multi-objective routing problem as
+where with .
+Here, represents a feasible route in the multigraph , encoded as a sequence of edges, and is the associated cost vector.
+Due to the combinatorial nature of the feasible solution space and the multiple conflicting objectives, the problem is generally NP-hard.
+The solution to an MO routing problem is a set of objective values known as the Pareto Front (PF) with associated routes in the Pareto Set (PS). Formally, we define the PS as
 where denotes that dominates , meaning
@@
 Intuitively, the PS represents solutions for which no strictly better alternative exists.
-To obtain this set, a typical method is decomposition through scalarization.
-The idea is to translate the MO problem into several SO problems that can be solved in parallel or consecutively.
+To obtain this set, we utilize a typical method: decomposition through scalarization.
+The idea is to translate the MO problem into several subproblems that can be solved separately.
 Let such that and be a vector that controls the preference among the objectives.
-Then, the most straight-forward scalarization is the Linear Scalarization (LS) .
-The LS subproblem is then , which is a standard SO problem.
-Unfortunately, it is well-known that there may be solutions in the Pareto Set that do not solve an LS subproblem for any preference (Ehrgott 2005).
+Then, under Linear Scalarization (LS), the single-objective cost is and the corresponding subproblem is .
+Unfortunately, it is well-known that there may be solutions in the PS that do not solve an LS subproblem for any preference (Ehrgott 2005).
 The Chebyshev scalarization has more favorable properties.
 Let be an ideal value for objective .
-Then the Chebyshev scalarized cost is
-Given a solution in the PS, there always exists a preference and corresponding subproblem for which the solution is optimal (Choo and Atkins 1983).
-In our context of multigraphs, it is interesting to consider how the problem simplifies under scalarization.
-We note that under LS, the MOTSP on a multigraph can be simplified to a simple graph by pruning edges that have sub-optimal LS cost compared to a parallel edge.
-Under Chebyshev scalarization or for more complicated problems this simple method will not work.
-Instead, we propose to learn which parallel edge is optimal while simultaneously learning to construct the optimal route.
+Then, the Chebyshev scalarized cost is
+Given a solution in the PS, there always exists a preference and corresponding Chebyshev subproblem for which the solution is optimal (Choo and Atkins 1983).
+In our context of multigraphs, it is interesting to consider how problems simplify under scalarization.
+We highlight an important result for the Multi-Objective TSP (MOTSP), in which the total cost of a tour is the sum of the vectorial costs of the traversed edges.
+The proof of the following proposition is omitted due to its simplicity.
+Given a linear scalarization defined by a preference vector , define the reduced graph by removing any edge whose scalarized cost is strictly worse than that of a parallel edge.
+Proposition 1.
+Let denote the linearly scalarized cost and let be the set of feasible tours on the pruned graph . Then, the optimal value of the scalarized subproblem is preserved:
+This illustrates that a multigraph representation is sometimes easy to handle using pruning.
+However, under Chebyshev scalarization or for more complex problems (e.g., with edge attributes that contribute nonlinearly to the cost), similar results might not be available.
+As such, we propose to learn which parallel edge is optimal while simultaneously learning to construct the optimal route for each preference .
+Moreover, by explicitly taking into account the multigraph representation when designing our models, we obtain better results and faster inference than if we pre-prune the multigraphs, even for the MOTSP.
 4 GNN-Based Multigraph Solver
-In the following, we detail two approaches based on Graph Neural Networks (GNNs) for solving MO multigraph routing problems.
-We call our method GNN-based Multigraph Solver (GMS) and propose an edge-based variant, which is slow but simple, as well as an architecture with two decoders, which is faster but more complex.
-Both of these utilize the Graph Edge Attention Network (GREAT) of Lischka et al.
-(2024) as the encoder, the reason being that it 1) Can handle input data structured as multigraphs; 2) Outputs edge-embeddings; 3) Has been applied in an autoregressive-framework and shown competitive performance for non-Euclidean SO problems.
-However, note that it is possible to use other architectures in the encoder, such as e.g., the anisotropic GNN of Sun and Yang (2023).
+In the following, we propose two approaches based on GNNs for solving MO multigraph routing problems.
+We call our method GNN-based Multigraph Solver (GMS) and propose an edge-based variant constructing routes autoregressively directly on the multigraph as well as a variant with two decoders that first prunes and then constructs routes.
+Both variants require an encoder that i) can handle input data structured as multigraphs and ii) works with and outputs edge embeddings.
+While there are multiple GNN architectures that fit this description, we utilize the Graph Edge Attention Network (GREAT) of Lischka et al.
+(2025), since it has been applied in an autoregressive framework and shown competitive performance for non-Euclidean single-objective routing problems.
 Specifically, we utilize the Node-Based (NB) version of GREAT.
@@
 4.1 Edge-Based GMS
-We visualize Edge-Based GMS (GMS-EB) in Figure 1.
+Vanilla node-based construction is insufficient for our purpose, as it relies on predicting the next node in the route, while we also require an edge-selection between the current node and the next.
+Thus, we propose an end-to-end edge-based model that instead predicts the next edge, and thereby implicitly the next node.
+We call this Edge-based GMS (GMS-EB).
+We visualize GMS-EB in Figure 1.
 The encoder, consisting of GREAT-layers, outputs edge embeddings.
-Using these, the decoder constructs valid tours auto-regressively.
-As in Lin, Yang, and Zhang (2022); Navon et al.
-(2021) we utilize a hyper-network to generate the weights of the decoder conditioned on the current preference.
-The decoder is an edge-based variant of the Multi-Pointer (MP) decoder from Jin et al.
-(2023), which we detail in Appendix A.
-Given the context of the current node in the tour, the decoder outputs probabilities of traversing edges starting in that node.
-Note that the underlying graphs in routing problems are usually complete.
-Hence there are edges in a graph of nodes.
-Moreover, the decoder needs to be run times to construct a full tour and we usually roll out sample trajectories.
-Hence an edge-based decoder scales according to compared to a node-based that scales according to .
-We find that this is a bottleneck in terms of memory requirements and inference time.
-Hence we propose a second architecture that utilizes a node-based decoder.
+Using them, the decoder constructs valid tours autoregressively.
+Given the instance and incomplete route in construction step , the decoder selects edge with probability .
+Thus the probability of the whole route is
+Here, is the weights of the model given the current preference .
+As Navon et al.
+(2021) and Lin, Yang, and Zhang (2022), we utilize a Multi-Layer Perceptron (MLP) hyper-network to preference-condition the decoder, while keeping the encoder preference-agnostic.
+Consequently we form the decoder weights according to and set
+Regarding the decoder architecture, it is an edge-based variant of the Multi-Pointer (MP) decoder from Jin et al. (2023). We detail this component further in Appendix B.
+Note that a multigraph with nodes and edges between each pair of nodes has edges in total.
+Moreover, the decoder must be run times to construct a full tour and we roll out sample trajectories according to the POMO framework (Kwon et al.
+2020).
+Hence, an edge-based decoder scales as compared to a node-based one that scales as .
+Consequently, we propose a second architecture that employs a node-based decoder and thereby scales better to higher node counts in terms of memory requirements and runtime.
 4.2 Dual Head GMS
-Using GREAT, we can obtain node-embeddings by not applying (2) in the final layer.
-The problem is that the action of selecting a node is not sufficient, as a path selection is also needed.
-Thus, we need to ensure that the action of choosing a node uniquely defines which edge to choose, which is achieved by pruning parallel edges leaving a unique one between each node-pair.
+For a node-based decoder, we need to ensure that the action of choosing a node uniquely defines which edge to choose.
+This is achieved by pruning parallel edges, leaving a unique one between each node pair.
 In general, choosing this edge is not trivial.
-Thus, we propose to learn the choice using a dual decoder setup.
+As such, we propose to learn the choice using a dual decoder setup.
 We call this approach Dual Head GMS (GMS-DH) and it is visualized in Figure 2.
-The key is that we insert a second decoder before the final encoder layer.
-This decoder uses the intermediate edge-embeddings to select which parallel edge to keep.
-The selection is fed back into the encoder, which only propagates the embedding of the remaining edges and converts them to node-embeddings.
-We also find it beneficial to add a small number of standard transformer layers after the final GREAT-layer in order to obtain more expressive node-embeddings.
-We detail the architecture for the selection-decoder in Appendix A.
-For the routing-decoder, we use a preference-conditioned version of the Multi-pointer decoder (Jin et al.
+The key is that we insert another decoder before the final GREAT layer, which uses intermediate edge embeddings to select the parallel edge to keep.
+Let be the set of edges between the node pair .
+The selection-decoder produces the probability of retaining .
+As we treat each node pair independently and select one edge per pair, the probability of the joint selection of active edges is
+Again, the weights are generated by an MLP hyper-network. We detail the architecture for the selection-decoder further in Appendix B.
+The edge choice from the first decoder is fed back into the encoder, where the embeddings of corresponding edges are aggregated into node embeddings using (7).
+After this final GREAT layer, we find it beneficial to add a small number of standard transformer layers to obtain more expressive node embeddings for the second decoder, which constructs routes.
+The routing-decoder is a preference-conditioned version of the multi-pointer decoder (Jin et al.
 2023).
-When applying GMS-DH to simple graphs, we simply remove the first decoder.
-Also note that this architecture ensures that the first layers only need to be run once for each instance, which reduces the run-time compared to selecting edges before encoding.
-From a broader perspective, our approach in GMS-DH can be viewed as a mixture of non-autoregressive and autoregressive learning.
-As a drawback, the edge-choice can no longer depend on the routing context, instead it must be chosen before the roll-out is started.
+Similarly to the edge-based version above, given the partial tour , this decoder outputs the probabilities for the next construction step, where now denotes a node and the probability depends on the edge choice .
+Remark that when applying GMS-DH to simple graphs, we simply remove the first decoder.
+Also note that this architecture ensures that the first layers only need to be run once for each instance as they are preference-agnostic and utilize the entire multigraph.
+This reduces the runtime compared to letting the encoder be preference-aware or pruning the multigraph for each preference before encoding.
+From a broader perspective, our approach in GMS-DH can be viewed as a combination of non-autoregressive and autoregressive construction.
+As a drawback, the edge choice can no longer depend on the context given by the partial tour.
+Instead, it must be chosen before starting roll-out.
 4.3 Training Algorithm
 Our first proposed model, GMS-EB, is trained using the Multi-Objective REINFORCE from Lin, Yang, and Zhang (2022).
-In short, given a preference and instance , the model parameterizes a policy .
-A tour is associated with a reward according to (1)
-The objective is to maximize
-Additionally, we form a baseline by rolling out tours for the same instance and averaging the reward, according to the POMO framework (Kwon et al.
-2020), to stabilize the training.
-The policy gradient is utilized to update the model weights and it is approximated, given one preference , according to
+Given a preference and instance , the model parameterizes a policy according to (9).
+A tour is associated with a Chebyshev reward corresponding to the negative scalarized cost (5):
+The objective during training is to maximize
+To stabilize the training, we form a baseline by rolling out tours for the same instance and averaging the reward, according to the POMO framework (Kwon et al.
+2020).
+The policy gradient is then utilized to update the model weights and it is approximated, given one preference , according to
 where is the batch size of instances and the advantage is
-For the second model, the training is more complicated.
-We can view each head as an agent.
-The selection-head takes actions corresponding to which edge to choose between each node pair.
-It parameterizes a policy .
-The routing-head uses the selection, which we denote by .
-It parameterizes a policy and obtains a scalar reward according to (3).
+For the second model, the training is more involved.
+Given an edge selection from the selection-head according to (11), the routing-head parameterizes a policy and obtains a scalar reward by (12) as usual.
 With a fixed , the objective for this head is to maximize
-On the other hand, the task of the selection-head is to ensure that the edge selection is optimal. Let denote the optimal tour given edge selection . Then, define
+On the other hand, the task of the selection-head is to ensure that the edge choice is optimal.
+Let denote the optimal tour on the pruned graph given selection .
+Then, define the reward and objective function based on the optimal tour according to
 As the optimal tour is not available, we approximate it using sampled tours :
@@
 Thus, we approximate the policy gradients with
-For easy reference, we outline the training of GMS-DH in Algorithm 1.
+We outline the training of GMS-DH in Algorithm 1.
 Input: Preference distribution , instance distribution , batch size , number of samples per instance head 1, , and head 2,
-In addition, for both models we utilize a curriculum learning approach to speed up training.
+In addition, we utilize a form of curriculum learning to speed up training for both models.
 We start by training on small graph sizes and then gradually increase the problem size.
 For GMS-DH in the multigraph setting, we also start by training the routing-head on simple graphs to ensure that the approximation (18) is accurate even during the initial epochs of multigraph training.
-Further details regarding the curriculum learning can be found in Appendix B.
+Further details regarding the curriculum learning can be found in Appendix C.
 5 Experimental Studies
@@
 5.1 Benchmarking Setup
-Problems: In the asymmetric simple graph case we consider two standard routing problems: The Multi-Objective Traveling Salesman Problem (MOTSP) and Multi-Objective Capacitated Vehicle Routing Problem (MOCVRP).
+Problems
+In the asymmetric simple graph case we consider two standard routing problems: The Multi-Objective Traveling Salesman Problem (MOTSP) and Multi-Objective Capacitated Vehicle Routing Problem (MOCVRP).
 Specifically, we consider bi-objective variants of these.
@@
 In the FLEX-distribution, we sample edge distances independently and remove edges which are dominated.
-In the FIX distribution we first sample edges independently, but then rearrange by sorting, so that the an edge that is best in one objective is the worst in the other objective and so on.
+In the FIX distribution we first sample edges independently, but then rearrange by sorting, so that the an edge that is best in one objective is the worst in the other objective.
 Thus, in the FIX-distribution there are always exactly edges between each node pair, whereas FLEX has at most edges.
-In Appendix C, we provide more in-depth explanations of the problems and distributions. We also provide results on Euclidean problems in Appendix E.
+In Appendix D, we provide more in-depth explanations of the problems and distributions. We also provide results on Euclidean problems in Appendix G.
 TMAT50 | TMAT100 | XASY50 | XASY100 | ||||||||||
@@
 OR Tools | 0.86 | 4.21% | (24m) | 0.91 | 3.46% | (1.8h) | 0.90 | 2.59% | (24m) | 0.93 | 2.25% | (2.0h) | |
-| MOEA/D | 0.74 | 17.50% | (18h) | 0.66 | 30.28% | (12h) | 0.78 | 15.42% | (18h) | 0.69 | 27.63% | (12h) | |
+MOEA/D | 0.74 | 17.50% | (18h) | 0.68 | 28.46% | (99h) | 0.78 | 15.42% | (18h) | 0.72 | 24.69% | (91h) | |
 MBM | 0.86 | 5.21% | (13s) | 0.91 | 3.72% | (44s) | 0.90 | 2.46% | (13s) | 0.93 | 2.41% | (44s) | |
@@
 GMS-DH (aug) | 0.89 | 1.45% | (1.5m) | 0.93 | 1.61% | (6.6m) | 0.91 | 1.40% | (1.5m) | 0.93 | 1.92% | (6.6m) | |
-Baselines: For the MOTSP and MGMOTSP, we utilize linear scalarization together with strong single-objective methods in the form of LKH3 (Helsgaun 2017) and Google OR-tools.
+GMS-DH PP | 0.77 | 14.24% | (33s) | 0.83 | 12.53% | (2.3m) | 0.82 | 10.51% | (33s) | 0.85 | 10.39% | (2.3m) |
+Baselines
+For the MOTSP and MGMOTSP, we utilize linear scalarization together with strong single-objective methods in the form of LKH3 (Helsgaun 2017) and Google OR-tools (Perron and Furnon 2019).
 We also benchmark against four Multi-Objective Evolutionary Algorithms (MOEAs): MOGLS (Ishibuchi and Murata 1998), MOEA/D (Zhang and Li 2007), NSGA-II (Deb et al.
@@
 Below, we only show the results for the MOEA that performs best in each scenario.
-Finally, we design a neural baseline using the MatNet-architecture (Kwon et al.
+Finally, we also design a neural baseline based on the MatNet-architecture (Kwon et al.
 2021) as encoder and a preference-conditioned MP network as decoder.
-We call this MatNet-Based Model (MBM) and our motivation is that this is a learning-based method that can be used almost ”off-the-shelf” by combining components from literature.
-In the multigraph setting, this architecture requires us to pre-process the input by pruning edges until we obtain a simple graph, which we do using linear scalarization.
-In Appendix A and D, we provide further details about the neural benchmark and MOEAs respectively. In Appendix E, we also show results for all MOEAs as well as some other weaker heuristics.
-Model Settings: We train all models for epochs using randomly generated instances per epoch with the ADAM optimizer (Kingma and Ba 2015) and a learning-rate of .
+We call this MatNet-Based Model (MBM) and our motivation is that it is a method that can be used almost ”off-the-shelf” by combining components from literature.
+In the multigraph setting, this architecture requires us to pre-process the input by pruning edges until we obtain a simple graph, which we do using linear scalarization following Proposition 1.
+Model Settings
+We train all models for epochs using randomly generated instances per epoch with the ADAM optimizer (Kingma and Ba 2015).
+We use the batch size and the learning-rate .
 For GMS-EB we let the number of layers be , while GMS-DH has and .
-For both models we use as the embedding dimension and attention heads in the GREAT layers and MP decoders.
+For both models we set the embedding dimension to 128 and use attention heads in the GREAT layers and MP decoders.
 Regarding the number of samples in the decoding, we set for the selection-decoder during training and during inference.
@@
 (2025).
-Metrics: We utilize the Hypervolume (HV) metric (Zitzler et al.
+Metrics
+We utilize the Hypervolume (HV) metric (Zitzler et al.
 2003) to evaluate the performance of the methods.
 Formally, for a given set of points and a reference point that is dominated by all points in , the HV is defined as:
-where denotes the Lebesgue measure in , and represents the axis-aligned hyperrectangle bounded by and . Intuitively, a higher HV generally indicates better performance.
-In our experiments, we report the normalized HV, which is scaled to lie in the interval , averaged over 200 test instances.
+where denotes the Lebesgue measure in , and represents the axis-aligned hyperrectangle bounded by and .
+Intuitively, a higher HV generally indicates better performance, since the points in the obtained pareto front are further from the reference.
+We report the normalized HV, which is scaled to lie in the interval , averaged over 200 test instances.
 For the neural models we use 101 preferences during inference, which is also the number of subproblems for the benchmarks relying on scalarization.
 Apart from the HV, we provide the HV gap compared to the best-performing method as well as the inference time.
-In Appendix B, we also report some training times for the learning-based methods.
-For these methods, we used one NVIDIA Tesla A40 GPU with 48GB VRAM during training and inference.
-The non-learning-based methods were all tested using an Intel Xeon Gold 6338 CPU and 8 parallel processes.
+In Appendix C, we also report some training times for the learning-based methods.
+Hardware
+We conducted training and inference for the learning-based methods using a single NVIDIA Tesla A40 GPU with 48GB of VRAM.
+For the non-learning-based methods, evaluations were performed on an Intel Xeon Gold 6338 CPU and, to enable a more fair comparison, we report execution time based on solving multiple instances in parallel.
+Akin to Kwon et al.
+(2021), we set the number of parallel processes to 8.
 5.2 Performance on Benchmark Problems
-We show the performance of GMS together with the benchmarks in Table 1.
-As can be seen, GMS outperforms all benchmarks except LKH on the MOTSP and MGMOTSP, while it is the best performing model on the MOCVRP.
+We show the performance of GMS together with the other methods in Table 1.
+Both variants of GMS outperform most benchmarks on all problems.
+On the MOCVRP they are the best performing methods, while only LKH outperform them on the MOTSP and MGMOTSP.
 This showcases that our architectures are generally applicable across a variety of distributions, instance sizes and problems.
-We note that the runtime for GMS-EB scales poorly with the problem size, but it is still fast compared to most benchmarks, especially without augmentation.
-| FLEX2-20 | FLEX2-50 | |||||
-| HV | Gap | Time | HV | Gap | Time | |
-| MOEA/D | 0.75 | 9.19% | (2.7h) | 0.60 | 32.79% | (34h) |
-| MBM | 0.72 | 12.89% | (4.9s) | 0.80 | 11.09% | (14s) |
-| MBM (aug) | 0.73 | 11.98% | (34s) | 0.81 | 10.10% | (1.8m) |
-| GMS-EB | 0.81 | 1.95% | (4.3s) | 0.89 | 0.97% | (60s) |
-| GMS-DH | 0.80 | 3.37% | (3.1s) | 0.85 | 5.10% | (13s) |
-| GMS-EB (aug) | 0.82 | 0.00% | (31s) | 0.90 | 0.00% | (7.9m) |
-| GMS-DH (aug) | 0.82 | 0.33% | (16s) | 0.87 | 2.69% | (1.6m) |
-5.3 Multigraph Problem without Clear Pruning
-The main motivation behind learning to select edges in the multigraphs is to plan routes when no clear pruning method exists.
-Thus, we evaluate the performance on such a problem as well.
-Specifically, we consider an MGMOTSP with Time-Windows (MGMOTSPTW).
-Each edge is associated with a time and distance, while the objectives correspond to the number of violated time-windows and the total distance.
-Unlike the MGMOTSP with linear scalarization, it is not clear in this case which edges can be pruned apriori due to the more intricate time-window objective.
-We display the performance on this problem in Table 2, and benchmark against MBM with LS pruning as well as MOEAs.
-Evidently, our models are the best performing ones, illustrating the value of learning to select edges in combination with learning-to-route.
+As for the baseline neural method, MBM, it performs quite well across many cases, especially on the MGMOTSP, but severely under performs on some distributions, e.g., XASY100 for the MOCVRP.
+However, MBM is usually slightly faster than GMS-DH, owing to the MatNet architecture being faster and more memory-efficient than GREAT (for a comparison, see the results by Lischka et al.
+(2025)).
+GMS-DH has a small advantage for the MGMOTSP with nodes due to only encoding each multigraph once, but this is offset again for nodes.
+Nonetheless, both GMS models remain efficient relative to the other benchmarks, despite the slower scaling of GMS-EB.
+In particular, GMS-DH as well as GMS-EB without augmentation are significantly faster than LKH.
+Apart from the previously mentioned benchmarks, we also evaluate a variant of GMS-DH with Pre-Pruning (GMS-DH PP) applied before the encoder in the MGMOTSP setting.
+This pre-pruning is the same as for MBM, and the selection head is removed accordingly.
+The poor performance of GMS-DH PP in Table 1 highlights the importance of explicitly designing for the multigraph structure, even when a theoretically sound pruning strategy is available.
+Since GMS-DH PP requires re-running the encoder for each preference, it is approximately 3 slower than GMS-DH.
+We also hypothesize that the pre-pruning step shifts the data distribution in a way that GMS-DH is ill-equipped to model, resulting in performance that is worse than MBM with the same pruning strategy.
+Indeed, in our experience GMS-DH PP tends to produce Pareto fronts with poor diversity, clustered around central preference regions
 5.3 Zero-Shot Generalization
-We also test our models in a zero-shot setting on larger instances, which is displayed in Table 3. Both GMS-DH and GMS-EB remain competitive across both problems and all distributions.
+We also test our models on larger instances not seen during training, and report the results in Table 2.
+Both GMS-DH and GMS-EB remain competitive across both problems and all distributions in this zero-shot setting.
 TMAT200 | XASY200 | ||||||
@@
 GMS-EB | 0.95 | 1.74% | (50m) | 0.96 | 1.34% | (50m) | |
-| GMS-DH | 0.95 | 2.27% | (2.2m) | 0.94 | 2.56% | (2.2m) | |
-5.5 Ablation Study
-Finally, we showcase the ability of the selection-decoder in GMS-DH to select beneficial edges.
-Table 4 shows the result when replacing this component with a simple pruning function that removes edges with suboptimal linearly scalarized cost.
-For fairness, we also benchmark using both linearly scalarized reward and Chebyshev reward.
-As discussed, in the linear case the pruning should be optimal.
-Evidently, Chebyshev reward improves the results substantially. Moreover, learning the selection compared to using a simple function also seems slightly better, even in the simple case of the MGMOTSP.
-| FLEX2-50 | ||
-| HV | Gap | |
-| GMS-DH (Ch. Reward) | 0.8862 | 1.84% |
-| GMS-DH (Lin. Reward) | 0.8781 | 2.74% |
-| GMS-DH Simple (Lin. Reward) | 0.8789 | 2.65% |
-| GMS-DH Simple (Ch. Reward) | 0.8847 | 2.00% |
-6 Conclusion and Future Work
-In this paper, we propose two methods to learn to construct solutions for routing problems on multigraphs.
-We highlight that while we solve problems with multiple objectives, our models can be valuable in all scenarios with no clear choice between alternative paths.
-In future works, we will apply our methods on more advanced problems.
-A key weakness of our approach is that GMS-DH does not quite reach the same performance as GMS-EB in most cases.
-In this regard, there are works that are tangential to ours that could be utilized to obtain even better results for either GMS-DH or GMS-EB.
-Examples include the HV training of Lu et al.
-(2024) and diversity enhancement of Chen et al.
-(2023b).
-In the future, we will also explore decomposition methods like Ye et al.
-(2024) to extend our approach to even larger problem sizes.
-Finally, from a broader perspective, we want to stress that routing problems on multigraphs are not well explored in literature.
-We believe one major reason is because of the complexity this formulation introduces.
-Learning-based methods can help alleviate this weakness and make the multigraph formulation more appealing.
+GMS-DH | 0.95 | 2.27% | (2.2m) | 0.94 | 2.56% | (2.2m) |
+5.4 More Difficult Multigraph Problem
+Finally, we revisit one of the main motivations behind this work: routing for problems without clear a priori edge selection.
+Thus, we evaluate the performance on a more complex problem.
+Specifically, inspired by Ben Ticha et al.
+(2017), we consider an MGMOTSP with Time-Windows (MGMOTSPTW).
+Each edge is associated with a time and distance, while the objectives correspond to the number of violated time-windows and the total distance.
+Remark that, unlike the MGMOTSP, it is not clear in this case which edges can be pruned due to the more intricate time-window objective.
+Besides MBM with linear pre-pruning and MOEAs, we benchmark against a simplified version of GMS-DH in order to ablate its ability to learn which edges are optimal.
+In this simplified version, we replace the selection-decoder with a simple function that removes edges with suboptimal linear cost.
+Compared to pre-pruning the multigraph before the encoder, this approach is smarter, as it removes the need to reencode.
+Furthermore, this simple variant actually performs quite well on the MGMOTSP, which we show and discuss in Appendix F.
+We display the results for MGMOTSPTW in Table 3.
+Our two models are the best performing ones by quite a large margin.
+Notably, vanilla GMS-DH also outperforms the simplified variant, which illustrates the value of learning to select beneficial edges in comparison to using a simple pruning heuristic.
+FLEX2-50 | FIX2-50 | |||||
+HV | Gap | Time | HV | Gap | Time | |
+MOEA/D | 0.60 | 32.79% | (34h) | 0.60 | 35.73% | (36h) |
+MBM | 0.80 | 11.09% | (14s) | 0.64 | 31.18% | (14s) |
+MBM (aug) | 0.81 | 10.10% | (1.8m) | 0.66 | 28.58% | (1.9m) |
+GMS-EB | 0.89 | 0.97% | (60s) | 0.93 | 0.55% | (59s) |
+GMS-DH | 0.85 | 5.10% | (13s) | 0.91 | 1.74% | (12s) |
+GMS-EB (aug) | 0.90 | 0.00% | (7.9m) | 0.93 | 0.00% | (7.9m) |
+GMS-DH (aug) | 0.87 | 2.69% | (1.6m) | 0.92 | 0.78% | (1.7m) |
+GMS-DH Simple | 0.83 | 7.01% | (13s) | 0.85 | 8.93% | (12s) |
+6 Conclusion
+In this paper, we propose two learning-based approaches for solving multi-objective routing problems, such as the MOTSP, on multigraphs and asymmetric graphs.
+Both methods utilize graph neural networks for encoding the graph structure but differ in their route construction strategies.
+The first method is edge-based and directly builds tours on the multigraph, while the second employs a two-stage decoding process, consisting of non-autoregressive pruning of the multigraph followed by route construction.
+Empirical results show that both methods perform well across a variety of settings.
+While the edge-based model is simpler and usually demonstrates slightly superior performance, the second model is more scalable due to lower memory and runtime demands.
+Finally, we highlight that our models can be valuable in all routing scenarios where multigraphs are appropriate, i.e., when there are multiple competing edges between each node pair.
+In future work, we plan to use our setup to address single-objective problems with hard constraints as well as stochastic variants.
 Acknowledgements
@@
 Through the affiliation of F.R., the work was also partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
-Appendix A Architecture Details
+Appendix A Extended Related Work
+In this section, we complement the literature review in the main text with a more thorough discussion of related work.
+A.1 Single-Objective Learning for Routing
+In the autoregressive paradigm, Vinyals, Fortunato, and Jaitly (2015) were early to propose learning solutions to the TSP.
+They utilized an oracle solver for supervised learning, and Bello et al.
+(2017) later introduced reinforcement learning for training.
+By using attention models and inherent symmetries in the problem, first Kool, van Hoof, and Welling (2018) and then Kwon et al.
+(2020) obtained greatly improved optimality gaps on various routing problems.
+Regarding asymmetric problems, Kwon et al.
+(2021) proposed MatNet, a mixed-attention architecture working on bi-partite graphs.
+More recent works are Jin et al.
+(2023); Drakulic et al.
+(2023); Chen et al.
+(2024); Bi et al.
+(2024).
+These papers have improved results, can handle constraints better and can scale to even larger problems than previously.
+Another recent notable contribution was made by Drakulic, Michel, and Andreoli (2025), who proposed a jointly trained generalist model to solve a multitude of combinatorial optimization problems, including examples from routing.
+In the non-autoregressive category, some works are Joshi, Laurent, and Bresson (2019); Fu, Qiu, and Zha (2021); Qiu, Sun, and Yang (2022); Sun and Yang (2023).
+Typically, these methods require more sophisticated procedures than simple sampling in the decoding, such as Monte–Carlo tree search, to perform competitively, but scale better to large instances due to lower space and time complexity.
+Thus, an interesting recent work combining autoregressive and non-autoregressive construction to solve large instances (up to nodes for the TSP) while maintaining performance is Ye et al.
+(2024).
+A.2 Multi-Objective Learning for Routing
+Early works treating MO routing using learning utilize multiple models, where one model is responsible for one preference between the objectives (Li, Zhang, and Wang 2021; Wu, Wang, and Zhang 2020).
+This quickly becomes infeasible as the number of models required grows exponentially with the number of objectives (Ehrgott 2005).
+Pareto-set learning for multi-objective routing using a single model was proposed by Lin, Yang, and Zhang (2022).
+They used the hyper-networks of Navon et al.
+(2021) to condition the decoder on the preference.
+Recent studies that feature improved results with a similar conditioning method include Fan et al.
+(2025) and Lu et al.
+(2024).
+An alternative approach is to use meta-learning to train a single model that can be quickly fine-tuned over a few steps to solve individual subproblems.
+Two representative works in this direction are Zhang et al.
+(2021) and Chen et al.
+(2023a).
+Another significant recent work is that of Chen et al.
+(2023b), in which two diversity enhancement methods are introduced.
+The authors combine these methods with both the meta-learning approach and hyper-network approach to achieve improved results.
+Besides tackling asymmetric problems, our work is novel in the MO routing setting as most methods utilize autoregressive construction, while one of our models combine autoregressive and non-autoregressive construction.
+A.3 Routing on Multigraphs
+Successful approaches to tackle MO shortest path problems on multigraphs include genetic algorithms (Beke, Weiszer, and Chen 2021), A∗-related methods (Weiszer, Burke, and Chen 2020) and hybrid approaches combining A∗ with learning (Liu, Chen, and Weiszer 2022).
+These methods all aim to find a Pareto set of shortest paths between two destinations.
+For combinatorial routing problems such as VRPs on multigraphs, heuristics and exact solvers have been proposed.
+The first study in this regard was conducted by Garaix et al.
+(2010), who designed a simple heuristic and branch-and-price method for a single-objective on-demand transportation problem.
+Other works for single-objective constrained problems or multi-objective problems are Lai, Caliskan Demirag, and Leung (2016); Ben Ticha et al.
+(2019); Tikani, Setak, and Demir (2021).
+In addition, Ben Ticha et al.
+(2017) clearly concluded that a multigraph setup yields significantly lower costs compared to a simple graph representation for the capacitated VRP with time-windows.
+At the same time, the computation time increases significantly under this perspective, and the authors highlight the need for new heuristics.
+Appendix B Architecture Details
 Here, we highlight some of the main components in the model architectures that are not described in the main text.
-A.1 Edge-Based Decoder in GMS-EB
-In GMS-EB, we utilize an edge-based multi-pointer decoder which is based on Jin et al.
+B.1 Edge-Based Decoder in GMS-EB
+In GMS-EB, we utilize an edge-based multi-pointer decoder which is inspired by Jin et al.
 (2023).
@@
 Finally, softmax is applied to transform the logits to probabilities.
-It must be noted that in our case, we preference-condition the decoder using a hyper-network. This is done by letting a network generate the weights and .
-A.2 Selection-Decoder in GMS-DH
+In our case, we also preference-condition the decoder using a hyper-network. This is done by letting a multi-layer perceptron generate the weights and .
+B.2 Selection-Decoder in GMS-DH
 The task of this decoder is to prune parallel edges in the intermediate stage of GMS-DH, leaving only one between each node pair.
@@
 In the experiments we set .
-These layers are not preference conditioned.
-Hence, they do not need to be rerun for each preference.
+These layers are not preference conditioned and hence they do not need to be rerun for each preference.
 Next, given two nodes, we want to compute the probability of retaining each (directed) edge in between.
@@
 Final edge probabilities are then computed using tanh-clipping and softmax as
-A.3 MatNet-Based Benchmark Model
+B.3 MatNet-Based Benchmark Model
 We visualize this architecture in Figure 3.
@@
 a distance matrix .
-To accommodate the use of several distance matrices, in our multi-objective setting, we follow the idea in Appendix A of the original paper.
+To accommodate the use of several distance matrices, , in our multi-objective setting, we follow the idea in Appendix A of the original paper.
 That is, we mix the internal attention score with all distance matrices using an MLP with inputs.
@@
 Thus, we must pre-prune a multigraph before passing it through the encoder.
-This can be done with some sort of heuristic in the general case, but, as discussed, under LS and considering the MOTSP, a simple pruning method is to remove edges with sub-optimal linearly scalarized cost.
+This can be done with some sort of heuristic in the general case, but under linear scalarization and considering the MOTSP, an optimal pruning method is to remove edges with suboptimal scalarized cost.
+We utilize this pruning method for both multigraph problems.
 Moreover, it can be noted that the encoder needs to be rerun for each new preference, which results in a slower architecture than if the encoding is done once.
-Appendix B Model Training
+Appendix C Model Training
 Here we outline the curriculum learning which we briefly discussed in the main text. We also present an analysis of training performance.
-B.1 Curriculum Learning Details
+C.1 Curriculum Learning Details
 For both GMS-EB and GMS-DH we utilize a simple curriculum learning approach, with the purpose to decrease training times and to improve performance on large problems.
@@
 This is because the selection-decoder relies on a reward signal from the routing-decoder.
-As such, we find that performance and rate of convergence improves if the routing-decoder is pre-trained to obtain a stable and accurate signal.
+As such, we find that performance and rate of convergence improves if the routing-decoder is pre-trained on simple graphs to obtain a stable and accurate signal.
 We also find that stability increases if we freeze the parameters of the shared encoder with respect to the loss of the selection-decoder.
 That is, only the loss from the routing-decoder is utilized to calculate encoder gradients.
-B.2 Training Times
+C.2 Training Times
 Table 4 shows the number of parameters for the models as well as training time per epoch for one problem on a Simple Graph (SG) and Multigraph (MG).
@@
 GMS-EB | 1.98M | 266K | 51m | 1.7h |
-Appendix C Routing Problems and Distributions
+Appendix D Routing Problems and Distributions
 In this section, we detail the distributions and routing problems which we solve.
-C.1 Distributions
+D.1 Distributions
 The following distributions for node distances are used across the problems on simple graphs.
@@
 Motivated by the fact that XASY seems more difficult to learn than TMAT, we hypothesize that the triangle inequality would contribute positively to model performance.
-C.2 MOTSP
-Problem Description: In the multi-objective traveling salesman problem the goal is to visit all nodes/customers and return to the starting position.
-The Euclidean version of this problem is widely treated in the learning-for-routing community, see e.g., Lin, Yang, and Zhang (2022) and Fan et al.
-(2025) in the multi-objective case.
+D.2 MOTSP
+Problem Description
+In the multi-objective traveling salesman problem the goal is to visit all nodes/customers and return to the starting position.
+The Euclidean version of this problem has been widely treated in the learning-for-routing community, e.g., by Lin, Yang, and Zhang (2022) and Fan et al.
+(2025).
 In our case, we sample two distances, corresponding to the two objectives, between each pair of customers.
@@
 A tour can start from any node.
-Data Generation: Edge distances are sampled according to the EUC, TMAT and XASY distributions above.
-Model Details: There are no node features for this problem, hence both the GREAT-based and MatNet-based encoders have the two distances as edge features, while the latter uses zero-vectors as ”row” input.
-Evaluation Details: For all distributions, we utilize , and as reference for HV calculation for , and nodes respectively.
-In the zero-shot experiments, we have as reference for 150 nodes and for 200 nodes.
-C.3 MOCVRP
-Problem Description: In the multi-objective capacitated vehicle routing problem, a truck starts at a depot and must pick up goods from a specified number of customers.
+Data Generation
+Edge distances are sampled according to the EUC, TMAT and XASY distributions above.
+Model Details
+There are no node features for this problem, hence both the GREAT-based and MatNet-based encoders have the two distances as edge features, while the latter uses zero-vectors as ”row” input.
+Evaluation Details
+For all distributions, we utilize , and as reference for HV calculation for , and nodes respectively. In the zero-shot experiments, we have as reference for 150 nodes and for 200 nodes.
+D.3 MOCVRP
+Problem Description
+In the multi-objective capacitated vehicle routing problem, a truck starts at a depot and must pick up goods from a specified number of customers.
 The vehicle has a fixed capacity while each customer has a demand.
@@
 As is common in the multi-objective setting (Lin, Yang, and Zhang 2022), we assume one objective is the traveled distance, while the other is the longest tour for a single vehicle (makespan).
-As such, each node-pair is only associated with one distance.
-Data Generation: We assume that the capacity of the truck is 30, 40 and 50 for 20, 50 and 100 nodes respectively, while the demands are sampled uniformly from the set .
+As such, each node pair is only associated with one distance.
+Data Generation
+We assume that the capacity of the truck is 30, 40 and 50 for 20, 50 and 100 nodes respectively, while the demands are sampled uniformly from the set .
 This is the same setting as in e.g., Kool, van Hoof, and Welling (2018) and Lin, Yang, and Zhang (2022).
 Edge distances are sampled according to the EUC, TMAT and XASY distributions.
-Model Details: In the decoding, we augment the routing context (represented by the query ) with the current load of the vehicle.
+Model Details
+In the decoding, we augment the routing context (represented by the query ) with the current load of the vehicle.
 For the GREAT-based encoders, we let the features of an edge be its distance as well as the demand of the customer at the end.
 On the other hand, the MatNet-based encoders only utilize the distance as an edge-feature while the demand is used as a node feature.
-Evaluation Details: For all distributions, we utilize , and as reference for HV calculation for the three considered instance sizes.
-C.4 MGMOTSP
-Problem Description: As for the MOTSP, the goal of the multigraph MOTSP is to visit all nodes and return to the start.
+Evaluation Details
+For all distributions, we utilize , and as reference for HV calculation for the three considered instance sizes.
+D.4 MGMOTSP
+Problem Description
+As for the MOTSP, the goal of the multigraph MOTSP is to visit all nodes and return to the start.
 The only difference is that the underlying structure is a multigraph.
 Similarly, we define two distances for each node pair and these are associated with the objectives of the problem.
-While this variant of the TSP on a multigraph has not been treated before to our knowledge (neither using learning nor other methods), some stochastic single-objective variants have been proposed in e.g., Tadei, Perboli, and Perfetti (2017) and Maggioni, Perboli, and Tadei (2014).
+While this variant of the TSP on a multigraph has not been treated before to our knowledge (neither using learning nor other methods), some stochastic single-objective variants have been proposed by e.g., Tadei, Perboli, and Perfetti (2017) and Maggioni, Perboli, and Tadei (2014).
 Moreover, we believe that the MGMOTSP is a suitable benchmark problem for our solvers, as it is simple and easy to find good baselines for.
-Data Generation: We sample distances between the nodes according to the FLEX and FIX distributions.
-Model Details: As the MOTSP, the MGMOTSP lacks edge features.
+Data Generation
+We sample distances between the nodes according to the FLEX and FIX distributions.
+Model Details
+As the MOTSP, the MGMOTSP lacks node features.
 Hence, we utilize the same setup for the models as in the MOTSP.
 Moreover, for the MatNet-based model, we sparsify the underlying multigraph given each preference before it is fed into the encoder.
-This is done by removing all edges with sub-optimal LS cost compared to a parallel edge.
-To make this pruning strategy exact, we train MBM using linearly scalarized reward instead of the Chebyshev reward.
-Evaluation Details: For the FLEX distribution, we utilize , and as reference for instance sizes 20, 50 and 100, regardless of .
+This is done by removing all edges with suboptimal linearly scalarized cost compared to a parallel edge.
+To make this pruning strategy exact, we train MBM using linear reward instead of the Chebyshev reward.
+Evaluation Details
+For the FLEX distribution, we utilize , and as reference for instance sizes 20, 50 and 100, regardless of .
 In the zero-shot experiments we set the references to and for FLEX with 150 and 200 nodes.
 For FIX, the reference is given by , where is the node count.
-C.5 MGMOTSPTW
-Problem Description: The goal in MGMOTSPTW is to create a tour that visits all customers and returns to the starting position.
+D.5 MGMOTSPTW
+Problem Description
+The goal in MGMOTSPTW is to create a tour that visits all customers and returns to the starting position.
 However, in this case we assume there is a fixed depot and that each customer/node is associated with a time-window.
@@
 2025), although not on multigraphs.
-In the multigraph framework, this problem is reminiscent of the vehicle routing problem with time-windows in Ben Ticha et al.
+In the multigraph framework, this problem is reminiscent of the vehicle routing problem with time-windows of Ben Ticha et al.
 (2017), but with a second objective that is the soft variant of the time-window constraint.
-Data Generation: Distances and travel times are sampled according to the FLEX and FIX distributions.
+Data Generation
+Distances and travel times are sampled according to the FLEX and FIX distributions.
 Additionally, we sample time-windows according to the ”medium” distribution of Chen et al.
@@
 We also argue that this premise is simple for initial evaluation of our models, especially in comparison to e.g., the ”hard” instances of Chen et al.
-(2024), which tests out-of-distribution generalization.
+(2024), which test out-of-distribution generalization.
 Further, we assume that the vehicle can depart immediately from each customer, i.e., that there is no service time, and that no waiting occurs, even if a vehicle arrives before the starting time.
@@
 We also argue that this assumption is fairly realistic.
-Model Details: In the decoder, we enhance the routing context with an embedding of the current time.
+Model Details
+In the decoder, we enhance the routing context with an embedding of the current time.
 In the encoder, the node features in this case are the starting time and end time of the service, while the edge features are distance and travel time.
@@
 For instance, removing an edge with a high travel time but low distance might not be optimal if traversing that edge would result in no violated time-windows anyway.
-Consequently, we apply the same simple pruning method as for the MGMOTSP.
-That is, we form the cost of each edge using a linear combination of its travel time and distance, then we remove all edges apart from the best one.
+For simplicity, we apply the same simple pruning method as for the MGMOTSP.
+That is, we form the scalarized cost of each edge using a linear combination of its travel time and distance, then we remove all edges apart from the best one.
 We also choose the linear reward instead of the Chebyshev reward for MBM.
-Evaluation Details: In this case, we let the reference value corresponding to the time-window violation be , while the reference value corresponding to the distance is the same as in the MGMOTSP case.
-Appendix D Setup for MOEAS
-As a base for the implementation of NSGA-II, NSGA-III and MOEA/D, we utilized Pymoo222https://pymoo.org/ and as a base for MOGLS, we utilized GeneticLocalSearch_TSP333https://github.com/kevin031060/Genetic˙Local˙Search˙TSP.
+Evaluation Details
+In this case, we let the reference value corresponding to the time-window violation be , while the reference value corresponding to the distance is the same as in the MGMOTSP case.
+Appendix E Setup for MOEAs
+As a base for the implementation of NSGA-II, NSGA-III and MOEA/D, we utilized Pymoo222https://pymoo.org/ (Blank and Deb 2020) and as a base for MOGLS, we utilized GeneticLocalSearch_TSP333https://github.com/kevin031060/Genetic˙Local˙Search˙TSP (Li, Zhang, and Wang 2021).
 For the former three, we limit the number of iterations to while the latter is limited to .
-Additionally, local search strategies is a necessity for efficient usage of genetic algorithms in combinatorial optimization.
+Additionally, local search strategies are a necessity for efficient usage of genetic algorithms in combinatorial optimization.
 For the MOTSP, we utilized the 2-opt heuristic of Jaszkiewicz (2002), while we utilized the problem-specific local search and chromosome decoding strategy of Lacomme, Prins, and Sevaux (2006) for the MOCVRP.
@@
 As this affects the runtime negatively (due to the need to search through all moves), we limit the number of local moves in each search depending on the problem size.
-Additionally, for the algorithms not requiring preference weights for each individual, we must calculate such weights to compare the improvements.
+Additionally, for the algorithms not requiring preference weights for each individual (NSGA-II and NSGA-III), we must calculate such weights to compare the improvements.
 In the bi-objective case, this is done using the strategy of Lacomme, Prins, and Sevaux (2006) according to
-where is the solution being evaluated and , are minimum and maximum values for the respective objectives.
+where is the solution being improved and , are minimum and maximum values for the respective objectives.
 In the multigraph setting, we design our own setup to tailor the MOEAs.
@@
 A local move is defined by a standard 2-opt move with the nodes, together with new edge selections given the changed node order.
-This edge selection is found by choosing the edge with the lowest LS cost between nodes that were previously not neighbors or have changed order.
-We want to stress that the focus in this study is not to design an evolutionary algorithms, we simply want a baseline to compare against.
+This edge selection is found by choosing the edge with the lowest linearly scalarized cost between nodes that were previously not neighbors or have changed order.
+We want to stress that the focus in this study is not to design evolutionary algorithms, we simply want baselines to compare against.
 We find that the proposed approach is somewhat efficient for small node counts, but performs quite badly for large graphs.
-Appendix E Complementary Numerical Results
+Appendix F Ablation Study
+Here, we present additional results obtained by replacing the selection-decoder in GMS-DH with a simple pruning function that discards edges with suboptimal linearly scalarized cost.
+As shown in Section 5.4, this approach performs rather poorly on the MGMOTSPTW.
+However, it is more reasonable to assume that it will yield good results for the MGMOTSP.
+To ensure a fair comparison, we evaluate both learned pruning and simple pruning using linearly scalarized rewards and Chebyshev rewards.
+In the linear case, simple pruning is optimal by Proposition 1.
+Moreover, the learned pruning model is trained using edge selections per instance.
+Accordingly, we use 4 more POMO samples when training the model with simple pruning to match the training budget.
+Ablation results are shown in Table 5.
+For the MGMOTSP, simple and learned pruning perform quite similarly (although learned pruning shows a slight advantage under Chebyshev rewards for both distributions).
+Indeed, these results illustrate that it can be a smart approach to replace the selection-decoder in GMS-DH if a sound heuristic or exact pruning method is available.
+However, pruning after the encoder clearly outperforms pre-pruning (as used in GMS-DH PP), both in terms of solution quality and runtime, since pre-pruning requires reencoding the graph for each preference.
+FLEX2-50 | FIX2-50 | |||
+HV | Gap | HV | Gap | |
+GMS-DH Learned (Ch. Reward) | 0.89 | 1.84% | 0.91 | 1.73% |
+GMS-DH Learned (Lin. Reward) | 0.88 | 2.74% | 0.91 | 1.66% |
+GMS-DH Simple (Ch. Reward) | 0.88 | 2.00% | 0.90 | 1.82% |
+GMS-DH Simple (Lin. Reward) | 0.88 | 2.65% | 0.91 | 1.76% |
+GMS-DH PP (Lin. Reward) | 0.77 | 14.24% | 0.82 | 10.51% |
+Appendix G Complementary Numerical Results
 In this section, we show additional numerical results for more distributions, benchmarks and instance sizes.
-E.1 Problems on Simple Graphs
+G.1 Problems on Simple Graphs
 Table 6 and 7 show results for the MOTSP and MOCVRP respectively.
-Apart from the benchmarks in the main text, we present the performance of three simple heuristics: Nearest Neighbor (NN), Nearest Insertion (NI) and Farthest Insertion (FI).
-As LKH and OR-tools, these utilize the linear scalarization to transform the MO problem into single-objective subproblems.
+Apart from the benchmarks in the main text, we present the performance of three simple greedy heuristics on the MOTSP: Nearest Neighbor (NN), Nearest Insertion (NI) and Farthest Insertion (FI).
+Like LKH and OR-tools, these utilize the linear scalarization to transform the MO problem into single-objective subproblems.
 Additionally, we show results on Euclidean instances (EUC).
@@
 This is to be expected though, as a common pattern in previous studies is that architectures built for directed graph inputs perform worse on EUC than those specialized for coordinate inputs (Lischka et al.
-2024; Kwon et al.
+2025; Kwon et al.
 2021).
@@
 GMS-DH (aug) | 0.60 | 1.29% | (10s) | 0.74 | 1.65% | (47s) | 0.79 | 1.21% | (4.7m) |
-E.2 Multigraph Problems
-We display additional results for the MGMOTSP and MGMOTSPTW in Tables 8-10.
+G.2 Multigraph Problems
+We display additional results for the MGMOTSP and MGMOTSPTW in Tables 8-11.
 As for the MOTSP, we include NN, NI and FI for MGMOTSP.
@@
 We call these methods implicit LKH and implicit OR-tools, as they implicitly minimize the time-window violations by reducing the time the tour takes.
+Additionally, for the MGMOTSP, we show results for FLEX and FIX, i.e., when there are 5 parallel edges between each node pair.
 FLEX2-20 | FLEX2-50 | FLEX2-100 | |||||||
@@
 FI | 0.75 | 11.98% | (12s) | 0.77 | 14.30% | (2.0m) | 0.83 | 12.20% | (14m) |
-| MOGLS | 0.71 | 16.29% | (1.9h) | 0.49 | 45.84% | (29h) | 0.43 | 54.91% | (18h) |
-| NSGA-II | 0.81 | 5.06% | (1.3h) | 0.74 | 17.62% | (19h) | 0.64 | 32.11% | (15h) |
-| NSGA-III | 0.80 | 5.82% | (1.3h) | 0.74 | 18.48% | (19h) | 0.63 | 33.16% | (14h) |
-| MOEA/D | 0.80 | 6.02% | (1.4h) | 0.74 | 17.50% | (18h) | 0.66 | 30.28% | (12h) |
+MOGLS | 0.71 | 16.29% | (1.9h) | 0.49 | 45.84% | (29h) | 0.34 | 64.38% | (129h) |
+NSGA-II | 0.81 | 5.06% | (1.3h) | 0.74 | 17.62% | (19h) | 0.64 | 31.98% | (103h) |
+NSGA-III | 0.80 | 5.82% | (1.3h) | 0.74 | 18.48% | (19h) | 0.63 | 33.01% | (104h) |
+MOEA/D | 0.80 | 6.02% | (1.4h) | 0.74 | 17.50% | (18h) | 0.68 | 28.46% | (99h) |
 MBM | 0.82 | 3.30% | (4.4s) | 0.86 | 5.21% | (13s) | 0.91 | 3.72% | (44s) |
@@
 GMS-DH (aug) | 0.84 | 0.86% | (19s) | 0.89 | 1.45% | (1.5m) | 0.93 | 1.61% | (6.6m) |
+FLEX5-20 | FLEX5-50 | FLEX5-100 | |||||||
+LKH | 0.93 | 0.00% | (1.3m) | 0.95 | 0.00% | (6.4m) | 0.97 | 0.00% | (30m) |
+OR Tools | 0.91 | 1.51% | (2.9m) | 0.93 | 2.01% | (25m) | 0.96 | 1.60% | (1.8h) |
+NN | 0.86 | 6.78% | (1.8s) | 0.90 | 5.24% | (7.2s) | 0.94 | 3.33% | (28s) |
+NI | 0.88 | 5.48% | (9.5s) | 0.89 | 6.41% | (1.7m) | 0.92 | 5.45% | (12m) |
+FI | 0.87 | 5.89% | (11s) | 0.89 | 7.03% | (2.0m) | 0.92 | 5.89% | (14m) |
+MOGLS | 0.84 | 9.71% | (2.4h) | 0.69 | 27.67% | (52h) | 0.57 | 41.21% | (222h) |
+NSGA-II | 0.88 | 5.19% | (1.4h) | 0.80 | 16.61% | (31h) | 0.71 | 26.79% | (215h) |
+NSGA-III | 0.87 | 5.85% | (1.4h) | 0.79 | 17.48% | (31h) | 0.70 | 27.98% | (215h) |
+MOEA/D | 0.88 | 5.07% | (1.5h) | 0.82 | 13.49% | (30h) | 0.76 | 22.05% | (210h) |
+MBM | 0.90 | 2.38% | (4.4s) | 0.93 | 2.48% | (13s) | 0.96 | 1.71% | (45s) |
+MBM (aug) | 0.92 | 1.26% | (32s) | 0.93 | 2.12% | (1.7m) | 0.96 | 1.44% | (5.9m) |
+GMS-EB | 0.92 | 0.87% | (8.1s) | 0.94 | 1.08% | (2.3m) | 0.96 | 1.11% | (23m) |
+GMS-DH | 0.92 | 1.03% | (3.5s) | 0.94 | 1.64% | (16s) | 0.96 | 1.66% | (1.3m) |
+GMS-EB (aug) | 0.92 | 0.59% | (1.1m) | 0.95 | 0.86% | (18m) | 0.96 | 0.96% | (3.0h) |
+GMS-DH (aug) | 0.92 | 0.64% | (23s) | 0.94 | 1.31% | (2.2m) | 0.96 | 1.45% | (9.9m) |
 FIX2-20 | FIX2-50 | FIX2-100 | |||||||
@@
 FI | 0.78 | 9.23% | (11s) | 0.84 | 8.82% | (2.0m) | 0.88 | 7.89% | (14m) |
-| MOGLS | 0.74 | 13.68% | (1.9h) | 0.65 | 29.64% | (29h) | 0.52 | 45.00% | (19h) |
-| NSGA-II | 0.81 | 5.64% | (1.3h) | 0.77 | 16.56% | (19h) | 0.69 | 27.94% | (15h) |
-| NSGA-III | 0.80 | 6.85% | (1.3h) | 0.75 | 18.79% | (19h) | 0.66 | 30.49% | (15h) |
-| MOEA/D | 0.80 | 6.20% | (1.4h) | 0.78 | 15.42% | (18h) | 0.69 | 27.63% | (12h) |
+MOGLS | 0.74 | 13.68% | (1.9h) | 0.65 | 29.64% | (29h) | 0.54 | 42.73% | (122h) |
+NSGA-II | 0.81 | 5.64% | (1.3h) | 0.77 | 16.56% | (19h) | 0.69 | 27.86% | (96h) |
+NSGA-III | 0.80 | 6.85% | (1.3h) | 0.75 | 18.79% | (19h) | 0.66 | 30.17% | (95h) |
+MOEA/D | 0.80 | 6.20% | (1.4h) | 0.78 | 15.42% | (18h) | 0.72 | 24.69% | (91h) |
 MBM | 0.84 | 1.57% | (4.4s) | 0.90 | 2.46% | (13s) | 0.93 | 2.41% | (44s) |
@@
 GMS-DH (aug) | 0.85 | 0.82% | (18s) | 0.91 | 1.40% | (1.5m) | 0.93 | 1.92% | (6.6m) |
+FIX5-20 | FIX5-50 | FIX5-100 | |||||||
+LKH | 0.84 | 0.00% | (56s) | 0.89 | 0.00% | (6.2m) | 0.91 | 0.00% | (29m) |
+OR Tools | 0.82 | 1.41% | (3.0m) | 0.87 | 1.78% | (25m) | 0.90 | 1.78% | (2.0h) |
+NN | 0.79 | 5.97% | (1.9s) | 0.85 | 4.22% | (7.2s) | 0.88 | 3.11% | (25s) |
+NI | 0.79 | 4.98% | 9.8s) | 0.84 | 5.27% | (1.7m) | 0.86 | 5.24% | (12m) |
+FI | 0.79 | 6.00% | (11s) | 0.83 | 6.45% | (2.0m) | 0.85 | 6.37% | (14m) |
+MOGLS | 0.74 | 11.00% | (2.5h) | 0.67 | 24.51% | (38h) | 0.62 | 32.11% | (253h) |
+NSGA-II | 0.79 | 5.96% | (1.5h) | 0.72 | 18.65% | (32h) | 0.63 | 30.45% | (216h) |
+NSGA-III | 0.77 | 7.78% | (1.5h) | 0.69 | 21.60% | (32h) | 0.60 | 33.77% | (218h) |
+MOEA/D | 0.79 | 5.44% | (1.5h) | 0.77 | 13.45% | (31h) | 0.71 | 22.40% | (230h) |
+MBM | 0.82 | 1.45% | (4.4s) | 0.87 | 2.34% | (13s) | 0.89 | 1.94% | (44s) |
+MBM (aug) | 0.83 | 0.47% | (32s) | 0.87 | 1.82% | (1.7m) | 0.90 | 1.69% | (5.9m) |
+GMS-EB | 0.83 | 1.05% | (8.1s) | 0.88 | 1.03% | (2.3m) | 0.90 | 1.08% | (23m) |
+GMS-DH | 0.82 | 1.75% | (3.5s) | 0.86 | 2.96% | (16s) | 0.88 | 3.44% | (1.3m) |
+GMS-EB (aug) | 0.83 | 0.77% | (1.1m) | 0.88 | 0.89% | (19m) | 0.90 | 0.99% | (3.0h) |
+GMS-DH (aug) | 0.83 | 1.21% | (23s) | 0.86 | 2.38% | (2.2m) | 0.89 | 2.98% | (9.8m) |
 FLEX2-20 | FLEX2-50 | |||||
@@
 GMS-DH (aug) | 0.82 | 0.33% | (20s) | 0.87 | 2.69% | (1.6m) |
-E.3 Generalization to Larger Instances
-In Tables 11-12, we report the generalization performance zero-shot for our methods in the MOTSP and MGMOTSP cases.
-It can be seen that both models retain competitive performance.
-In particular, GMS-EB often performs close to LKH while being faster.
+GMS-DH Simple | 0.76 | 8.18% | (3.3s) | 0.83 | 7.01% | (13s) |
+FIX2-20 | FIX2-50 | |||||
+HV | Gap | Time | HV | Gap | Time | |
+Implicit LKH | 0.28 | 67.24% | (1.3m) | 0.15 | 84.24% | (6.5m) |
+Implicit OR Tools | 0.25 | 71.02% | (3.4m) | 0.14 | 85.49% | (26m) |
+MOGLS | 0.62 | 28.85% | (2.8h) | 0.41 | 55.72% | (21h) |
+NSGA-II | 0.73 | 15.93% | (2.7h) | 0.56 | 39.66% | (35h) |
+NSGA-III | 0.71 | 18.88% | (2.5h) | 0.53 | 43.42% | (35h) |
+MOEA/D | 0.74 | 14.99% | (2.8h) | 0.60 | 35.73% | (36h) |
+MBM | 0.53 | 38.64% | (4.9s) | 0.64 | 31.18% | (14s) |
+MBM (aug) | 0.54 | 38.10% | (35s) | 0.66 | 28.58% | (1.9m) |
+GMS-EB | 0.86 | 1.17% | (4.3s) | 0.93 | 0.55% | (59s) |
+GMS-DH | 0.85 | 2.47% | (3.4s) | 0.91 | 1.74% | (12s) |
+GMS-EB (aug) | 0.87 | 0.00% | (31s) | 0.93 | 0.00% | (7.9m) |
+GMS-DH (aug) | 0.87 | 0.46% | (20s) | 0.92 | 0.78% | (1.7m) |
+GMS-DH Simple | 0.76 | 12.07% | (3.2s) | 0.85 | 8.93% | (12s) |
+G.3 Generalization to Larger Instances
+In Tables 12-14, we report more generalization results for our methods in the MOTSP and MGMOTSP cases.
+Notably, even when introducing significantly more edges by testing on FLEX and FIX, our methods remain competitive zero-shot.
 EUC150 | EUC200 | |||||
@@
 GMS-DH ZS | 0.88 | 5.10% | (30s) | 0.90 | 5.22% | (1.0m) |
+FLEX10-100 | FIX10-100 | |||||
+HV | Gap | Time | HV | Gap | Time | |
+LKH | 0.98 | 0.00% | (16m) | 0.86 | 0.00% | (15m) |
+OR Tools | 0.98 | 0.87% | (57m) | 0.84 | 1.63% | (1.1h) |
+GMS-EB ZS | 0.98 | 0.90% | (23m) | 0.84 | 1.49% | (23m) |
+GMS-DH ZS | 0.97 | 1.53% | (1.1m) | 0.82 | 4.11% | (1.1m) |
 FLEX2-150 | FLEX2-200 | |||||
@@
 GMS-DH ZS | 0.94 | 2.17% | (1.0m) | 0.95 | 2.27% | (2.2m) |
+FLEX5-150 | FLEX5-200 | |||||
+LKH | 0.98 | 0.00% | (39m) | 0.98 | 0.00% | (1.2h) |
+OR Tools | 0.97 | 1.35% | (2.1h) | 0.97 | 1.15% | (4.1h) |
+GMS-EB ZS | 0.97 | 1.05% | (45m) | 0.97 | 1.02% | (2.1h) |
+GMS-DH ZS | 0.96 | 1.86% | (1.7m) | 0.96 | 2.03% | (3.5m) |
 FIX2-150 | FIX2-200 | |||||
@@
 GMS-DH ZS | 0.94 | 2.38% | (1.1m) | 0.94 | 2.56% | (2.2m) |
+FIX5-150 | FIX5-200 | |||||
+LKH | 0.92 | 0.00% | (37m) | 0.93 | 0.00% | (1.2h) |
+OR Tools | 0.91 | 1.69% | (2.5h) | 0.92 | 1.59% | (4.5h) |
+GMS-EB ZS | 0.92 | 1.02% | (46m) | 0.92 | 0.98% | (2.1h) |
+GMS-DH ZS | 0.89 | 3.99% | (1.7m) | 0.89 | 4.46% | (3.5m) |